### v1

# A Class of De Giorgi Type and Hölder Continuity for Some Problems in Musielak-Orlicz-Sobolev Spaces

Abstract.
In this paper, we introduce a new class of De Giorgi type functions, denoted by $\mathcal{B}_{G(x,t)}$, and establish the Hölder continuity of its elements under suitable additional assumptions on the generalized N-function $G(x,t)$. As an application, we prove the Hölder continuity of solutions to quasilinear equations whose principal part is in divergence form with $G(x,t)$-growth conditions, including both critical and standard growth cases. The novelty of our work lies in the generalization of the Hölder continuity results previously known for variable exponent [30, X, Fan and D. Zhao] and Orlicz [51, G. M. Lieberman] problems. Moreover, our results encompass a wide variety of quasilinear equations.
Keywords. Class of De Giorgi Type; Hölder Continuity; Musielak-Orlicz-Sobolev Spaces; Weak solutions.
2020 Mathematics Subject Classification. Primary: 35R09, 35B65; Secondary: 47G20, 35K55
1. Introduction
1.1. Overview
The qualitative theory of partial differential equations studies the behavior of weak solutions $u\in W^{1,1}_{\mathrm{loc}}(\Omega)$ of problems of the form
$\operatorname{div}A(x,u,\nabla u)=B(x,u,\nabla u)\quad\text{in }\Omega.$ (1.1)
Existence results, based on the calculus of variations and topological methods, usually give solutions only in Sobolev spaces. To use these solutions in applications, it is important to know if they are more regular. In particular, one wants Hölder continuity $u\in C^{0,\alpha}_{\mathrm{loc}}(\Omega)$. This property lies between boundedness ($L^{\infty}$) and differentiability ($C^{1,\alpha}$), and is crucial for uniqueness, stability, and free boundary problems.
The study of regularity began with the famous works of De Giorgi [21] and Nash [73], later completed by Moser [70, 71]. They proved that weak solutions of linear elliptic equations in divergence form,
$\operatorname{div}(A(x)\nabla u)=0,$
with bounded measurable coefficients $A(x)$ satisfying uniform ellipticity $\lambda|\xi|^{2}\leq\langle A(x)\xi,\xi\rangle\leq\Lambda|\xi|^{2}$, are Hölder continuous. This result was very surprising, because it did not require smoothness of $A(x)$. It introduced two powerful techniques, the Moser iteration and the De Giorgi measure argument. Later, attention moved to quasilinear equations, such as the $p$-Laplacian
$\Delta_{p}u:=\operatorname{div}(|\nabla u|^{p-2}\nabla u).$
Uraltseva [80], Uhlenbeck [79], and Lieberman [50] proved that solutions of $\Delta_{p}u=0$ are $C^{1,\alpha}$. But this regularity is delicate for $\Delta_{p}u=f$, if $f\in L^{q}$ with $q<n/p$, the solutions may not even be continuous [35, 56, 55]. This shows that the regularity program must proceed in two steps: first, prove $u\in L^{\infty}$ (by De Giorgi’s lemma or Moser iteration), then prove $u\in C^{0,\alpha}$.
These ideas and techniques were systematically extended to tackle the more general non-polynomial growth of operators in the Orlicz setting. The foundational work by Donaldson and Trudinger [25] first established the boundedness ($L^{\infty}$) of solutions. Lieberman [51], who proved that weak solutions to quasilinear equations of the form $\operatorname{div}(a(|\nabla u|)\nabla u)=0$ are locally Hölder continuous ($C^{0,\alpha}$). This required the associated N-function $G(t)$ to satisfy the condition $\Delta_{2}$, which is essential for the Sobolev-Poincaré inequalities and the scaling arguments of the De Giorgi-Nash-Moser theory. These contributions laid the foundation of the regularity theory in the Orlicz framework.
Although Orlicz spaces encompass many classical examples, there remain important cases that they do not cover. For this reason, considerable research has focused on the variable exponent framework, where the growth is governed by a function $p(x)$. The analysis of problems involving terms of the type $t^{p(x)}$, initiated by Fan and Zhao [30], revealed a new regularity phenomenon. They proved that weak solutions of equations such as
$\operatorname{div}(|\nabla u|^{p(x)-2}\nabla u)=0,$ are locally Hölder continuous, provided the exponent function $p(x)$ is log-Hölder continuous. This condition, expressed as
$|p(x)-p(y)|\leq\frac{C}{-\ln|x-y|}\quad\text{for }|x-y|<1,$
was shown to be not only sufficient but essentially necessary for regularity, a fact later solidified by the work of Acerbi and Mingione [3].
A further major advancement in the theory of non-standard growth problems came with the analysis of double-phase functionals, which introduce a new type of heterogeneity: the growth exponent remains fixed, but the structure of the operator changes depending on the location in the domain. For functionals of the form
$\operatorname{div}\!\left(|\nabla u|^{p-2}\nabla u+a(x)|\nabla u|^{q-2}\nabla u\right),$
with $1<p\leq q$, the regularity theory reveals a novel feature: the regularity of solutions depends directly on the regularity of the coefficient $a(x)$. Colombo and Mingione [15, 16] proved that in the natural growth case, boundedness and Hölder continuity of minimizers and solutions hold provided $a(x)$ is Hölder continuous, i.e., $a(x)\in C^{0,\alpha}$. This condition is sharp: if $a(x)$ is merely continuous, minimizers may fail to be locally bounded [15]. This phenomenon highlights a key principle: in problems with non-uniformly elliptic structures, the regularity of the coefficients controlling the growth is often necessary for the regularity of solutions. The boundedness result in this setting was also independently obtained by Baroni, Colombo, and Mingione [7].
Due to the vast scope and rapid development of regularity theory, we cannot cite all contributions individually. We acknowledge with gratitude the many valuable works not mentioned here and instead refer the reader to comprehensive surveys and bibliographies, such as [57, 58]. For detailed discussions of specific regularity results, we direct the reader to the following foundational works and the extensive literature contained therein: for the classical Laplacian and linear theory, see [21, 73, 70, 71]; for the $p$-Laplacian and $(p,q)$-growth, see [39, 38, 8, 37, 80, 45, 48, 50, 79, 9, 10, 46, 19, 47, 52, 53, 56, 54, 1, 2, 32, 36, 33, 27, 42, 78]; for Orlicz and other non-standard growth conditions, see [51, 40, 18, 26, 12, 11, 59]; for double-phase problems, see [15, 16, 7, 20]; and for variable exponent $p(x)$-Laplacian problems, see [30, 31, 3, 17, 43].
The Musielak-Orlicz growth framework represents the most general setting, encompassing all the previously mentioned cases: standard $p$-growth, Orlicz, variable exponent, and double-phase (or $(p,q)$) growth. To the best of our knowledge, the literature contains only one other work specifically devoted to the Hölder continuity of solutions to divergence-form problems under such general Musielak-Orlicz growth conditions, namely that of Wang, Liu, and Zhao [81]. In their analysis, the authors impose a key structural assumption on the function $G(x,t)$:
There exists a strictly increasing differentiable function $\xi:[0,+\infty)\to[0,+\infty)$ that satisfies
$n\xi(s)>s\xi^{\prime}(s)>\xi(s)\quad\text{for all }s\geq 0,$
such that:
  • (A11)
    $G(x,st)\geq\xi(s)G(x,t)$ for all $s\geq 0$, $t\in\mathbb{R}$, and $x\in\Omega$;
  • (A12)
    $\xi,\xi^{-1},\xi_{\ast},\widehat{\widetilde{\xi}}\in\Delta_{\mathbb{R}^{+}}$.
A significant limitation of this condition is its restricted applicability. The assumption (A11), in particular, is not satisfied by many fundamental models central to the theory of non-standard growth problems. Notably, it fails to hold for variable exponent spaces, $(p,q)$-growth equations, and double-phase functionals, precisely the paradigmatic cases that a comprehensive Musielak-Orlicz theory should seek to unify. Consequently, while providing an interesting result under a specific set of constraints, the framework of [81] does not offer the generality required to recover these critical classical and modern examples.
Motivated by the preceding discussion, in this paper we establish the local Hölder continuity of weak solutions to (1.1) under general Musielak–Orlicz growth conditions on the operators $A$ and $B$. The principal novelty of our work lies in providing a unified framework that simultaneously covers the classical cases of the $p$-Laplacian, variable exponent, and Orlicz-type equations, while also treating both subcritical and natural growth (critical) cases with either Dirichlet or Neumann boundary conditions. Note that equation (1.1) involves a critical term, which makes the study of regularity more delicate and necessitates a careful analysis of the associated Sobolev conjugate for generalized Young functions. The study of such conjugates was initiated by Fan [28], who introduced the concept under the assumption that $A$ satisfies a Lipschitz condition in the spatial variable. This theory was significantly advanced by Cianchi and Diening [13], who established a sharper formulation under weaker regularity assumptions. However, despite these advances, a complete theory of Sobolev conjugates for generalized Young functions remains incomplete, lacking crucial structural properties such as the $\Delta_{2}$-condition, which is notoriously difficult to verify.
As explained above, our work has two main contributions. First, we establish the local Hölder continuity of weak solutions to (1.1) under general Musielak–Orlicz growth conditions on the operators $A$, which in particular generalizes the variable-exponent type operators treated in [30]. Moreover, our work closes the gap left in [30], where the critical cases were excluded. At the same time, it extends another class of equations in the framework of $(p,g)$-growth with critical term; see Subsection 1.3.
Four main difficulties arise from the Musielak–Orlicz structure when establishing regularity. First, the energy is nonhomogeneous and exhibits explicit $x$-dependence alongside the natural $t$-dependence, which prevents the direct application of classical scaling and iteration techniques such as the De Giorgi method. Second, the regularity proofs are highly sensitive to the precise parameters controlling the growth and structure of the function $G(x,t)$, requiring careful balancing of assumptions to ensure the validity of key inequalities. Third, even when adapted versions of De Giorgi-type methods are applicable, they may not guarantee regularity, as demonstrated by the counterexamples of Giaquinta [35] and Marcellini [54, 56], which show that weak solutions may fail to be regular even under standard structural assumptions. Finally, the critical case introduces additional challenges compared to the subcritical case, due to the lack of uniform integrability and the need for more refined estimates. These aspects collectively indicate that establishing regularity under non-standard growth conditions requires new ideas and a delicate analysis beyond the classical framework.
Our approach is motivated by the works of Lieberman [51] and Fan–Zhao [31], and in particular by the classical techniques of De Giorgi [21] and Ladyzhenskaya–Ural’tseva [49]. These foundational contributions introduced powerful tools for proving Hölder continuity of solutions. In particular, they defined the class $\mathcal{B}(\Omega,M,\gamma,\gamma_{1},\delta,1/q)$ and showed that all functions in this class are Hölder-continuous. This framework is especially effective under standard $m$-growth conditions (see, e.g. [34, 36, 49, 69]). For the treatment of critical cases, we further rely on the method developed by Ho, Kim, Winkert, and Zhang [43], and we also make use of some fundamental properties of the Sobolev conjugate established in [6].
1.2. Assumptions and Main Results
For the sake of clarity, in this subsection, we present only the main assumptions, key results, and core definitions. Additional definitions and technical details will be introduced in the subsequent sections.
Let $\Omega$ be a bounded domain in $\mathbb{R}^{n}$ with $n\geq 2$, and let
$G(x,t)=\int_{0}^{|t|}g(x,s)\,\mathrm{d}s,$
be a generalized N-function (see Section 2.2 for the precise definitions) satisfying the following structural conditions:
There exist constants $1<g^{-}\leq g^{+}<n$ such that for all $x\in\mathbb{R}^{n}$, $t>0$:
$1<g^{-}\leq\frac{g(x,t)t}{G(x,t)}\leq g^{+}.$ (G0)
There exists $F\geq 1$ such that for almost every $x\in\mathbb{R}^{n}$:
$F^{-1}\leq G(x,1)\leq F.$ (G1)
There exists $\mu\in(0,1]$ such that for every ball $B\subset\mathbb{R}^{n}$ with $|B|\leq 1$, for every $t\in[1,1/|B|]$, and for almost all $x,y\in B$:
$\mu G^{-1}(x,t)\leq G^{-1}(y,t).$ (G2)
For every $x_{0}\in\Omega$ and all sufficiently small $R>0$, there exists $L_{0}>1$ such that for all $t>1$:
$\eta(R):=\sup_{x,y\in B_{R}(x_{0})}\left|r\left(x,t\right)-r\left(y,t\right)\right|\leq\frac{L_{0}}{|\ln(2R)|},$ (G3)
where $r(x,t):=\dfrac{tg(x,t)}{G(x,t)}$.
We now introduce a De Giorgi-type class of functions tailored to our framework:
Definition 1.1.
Let $M,\gamma,\gamma_{1}$, and $\delta$ be positive constants with $\delta\leq 2$. A function $u$ in the Musielak-Orlicz-Sobolev space $W^{1,G(x,t)}(\Omega)$ (The definition of Musielak–Sobolev spaces is provided in Subsection 2.2), satisfying $\max_{\Omega}|u(x)|\leq M$, is said to belong to the class $\mathcal{B}_{G(x,t)}(\Omega,M,\gamma,\gamma_{1},\delta)$ if the functions $w(x):=\pm u(x)$ satisfy the inequality
$\int_{A_{k,\rho-\sigma\rho}}G(x,|Dw|)\,\mathrm{d}x\leq\gamma\int_{A_{k,\rho}}G\left(x,\left|\frac{w(x)-k}{\sigma\rho}\right|\right)\,\mathrm{d}x+\gamma_{1}|A_{k,\rho}|,$ (1.2)
for every ball $B_{\rho}\subset\Omega$, every $\sigma\in(0,1)$, and every threshold $k$ satisfying
$k\geq\max_{B_{\rho}}w(x)-\delta M,$ (1.3)
where $A_{k,\rho}:=\{x\in B_{\rho}:w(x)>k\}$.
One of the central results of this work is the following theorem:
Theorem 1.2.
Suppose that the function $G(x,t)$ satisfies conditions (G0)–(G1) and (G3). Then the De Giorgi-type class belongs to a Hölder space:
$\mathcal{B}_{G(x,t)}(\Omega,M,\gamma,\gamma_{1},\delta)\subset C^{0,\alpha}(\Omega),$
where the Hölder exponent $\alpha\in(0,1]$ depends only on the parameters $n,g^{-},g^{+},F,L:=\exp{L_{0}},\gamma$, and $\delta$, and is independent of $M$ and $\gamma_{1}$.
Definition 1.3 (see Sect. 1, Chapter 1 [49]).
We will say that $u\in\mathcal{B}_{G(x,t)}(\overline{\Omega},M,\gamma,\gamma_{1},\delta)$ if $u\in\mathcal{B}_{G(x,t)}(\Omega,M,\gamma,\gamma_{1},\delta)$ and the inequalities (1.2) hold also for arbitrary ball $B_{\rho}$ with $B_{\rho}\cap\partial\Omega\neq\phi$, $\sigma\in(0,1)$ and
$k\geq\max\left\{\max_{\Omega_{\rho}}w(x)-\delta M,\max_{\partial\Omega_{\rho}}w(x)\right\},$
where $\Omega_{\rho}=B_{\rho}\cap\Omega$, $\partial\Omega_{\rho}=\partial\Omega\cap B_{\rho}$, $A_{k,\rho}=\{x\in\Omega_{\rho}:w(x)>k\}$.
Definition 1.4 (see Sect. 1, of Chap. I of [49]).
We will say that the boundary $\partial\Omega$ of $\Omega$ satisfies the condition (A), if there exist two positive constants $m_{0}$ and $\theta_{0}$ such that for any ball $B_{\rho}$ with center on $\partial\Omega$ and radius $\rho\leq m_{0}$ and for any connected branch $\Omega^{\prime}_{\rho}$ of $B_{\rho}\cap\Omega$, the following inequality holds:
$|\Omega^{\prime}_{\rho}|\leq(1-\theta_{0})|B_{\rho}|.$ (A)
Our second main result is the following theorem.
Theorem 1.5.
Let $u\in\mathcal{B}_{G(x,t)}(\overline{\Omega},\gamma,\gamma_{1},\delta)$. If $G(x,t)$ satisfies conditions (G0)–(G1) and (G3), $\partial\Omega$ satisfies the condition (A) and $u|_{\partial\Omega}\in C^{0,\alpha_{1}}(\partial\Omega)$, then $u\in C^{0,\alpha}(\overline{\Omega})$ where $\alpha=\alpha(n,g^{-},g^{+},F,L,\delta,\alpha_{1})$.
As an application of Theorems 1.2 and 1.5, we investigate the Hölder continuity of weak solutions to the elliptic equation in divergence form:
$\operatorname{div}A(x,u,Du)+B(x,u,Du)=0,\quad x\in\Omega,$ ($\mathcal{P}$)
with Dirichlet boundary condition:
$\left\{\begin{array}[]{rcl}\operatorname{div}A(x,u,Du)+B(x,u,Du)&=&0,\quad x\in\Omega,\\ u&=&0,\quad\text{on }\partial\Omega,\end{array}\right$. ($\mathcal{PD}$)
or Neumann boundary condition:
$\left\{\begin{array}[]{rcl}\operatorname{div}A(x,u,Du)+B(x,u,Du)&=&0,\quad x\in\Omega,\\ A(x,u,Du)\cdot\nu&=&C(x,u),\quad\text{on }\partial\Omega,\end{array}\right$. ($\mathcal{PN}$)
where $\nu(x)$ denotes the outward unit normal vector to $\partial\Omega$ at $x\in\partial\Omega$. The functions
$A:\Omega\times\mathbb{R}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n},\quad B:\Omega\times\mathbb{R}\times\mathbb{R}^{n}\rightarrow\mathbb{R},\quad C:\partial\Omega\times\mathbb{R}\rightarrow\mathbb{R}$
are Carathéodory functions satisfying the following $G(x,t)$-structure conditions:
$A(x,u,\eta)\cdot\eta$ $\geq a_{4}\,G(x,|\eta|)-a_{5}\,H_{1}(x,|u|)-a_{6},$ ($\mathcal{A}_{1}$)
$|A(x,u,\eta)|$ $\leq a_{1}\,g(x,|\eta|)+a_{2}\,P(x,|u|)+a_{3},$ ($\mathcal{A}_{2}$)
$|B(x,u,\eta)|$ $\leq b_{1}\,M(x,|\eta|)+b_{2}\,h_{1}(x,|u|)+b_{3},$ ($\mathcal{B}$)
$|C(x,u)|$ $\leq c_{1}\,h_{2}(x,|u|)+c_{2}.$ ($\mathcal{C}$)
The auxiliary functions $M(x,t)$ and $P(x,t)$ are defined as:
$M(x,t)$ $=\widetilde{H_{1}}^{-1}(x,G(x,t)),$ (1.4)
$P(x,t)$ $=\widetilde{G}^{-1}(x,H_{1}(x,t)).$ (1.5)
The constants $a_{i}$, $b_{j}$, and $c_{k}$ are all positive, and we fix parameters $F_{1}>1$ and $F_{2}>1$. We also introduce the generalized N-functions:
$H_{1}(x,t)=\int_{0}^{t}h_{1}(x,s)\,\mathrm{d}s,\qquad H_{2}(x,t)=\int_{0}^{t}h_{2}(x,s)\,\mathrm{d}s,$
which satisfy the growth conditions:
$F_{i}^{-1}\leq H_{i}(x,1)\leq F_{i},$ $\qquad\text{for all }x\in\Omega,\quad i=1,2,$ ($\mathcal{H}_{1}$)
$g^{+}\leq h_{i}^{-}\leq\frac{h_{i}(x,t)\,t}{H_{i}(x,t)}\leq h_{i}^{+},$ $\qquad\text{for all }x\in\Omega,\ t>0,\quad i=1,2,$ ($\mathcal{H}_{2}$)
$\int_{\Omega}H_{i}(x,t)\,\mathrm{d}x<\infty,\ \text{for }t>0,\quad i=1,2,$ ($\mathcal{H}_{3}$)
and
$G(x,t)\prec H_{i}(x,t)\prec\prec G^{*}(x,t),\quad i=1,2,$ ($\mathcal{H}_{4}$)
where the symbol $\prec\prec"$ will be introduced in Definition 2.5, and
$G^{*}(x,t)=\int_{0}^{t}g^{*}(x,s)\,\mathrm{d}s$
is the Sobolev conjugate of $G(x,t)$ (see Subsection 2.2).
Remark 1.6.
The generalized N-functions $H_{1}(x,t)$ and $H_{2}(x,t)$ may exhibit critical growth corresponding to the conjugate $G^{*}(x,t)$. In such a case, we have $h_{i}^{+}=g_{*}^{+}:=\frac{ng^{+}}{n-g^{+}}$ and $h_{i}^{-}=g_{*}^{-}:=\frac{ng^{-}}{n-g^{-}}$ for $i=1,2$.
Definition 1.7.
Under assumptions (G0)–(G2), ($\mathcal{A}_{1}$)–($\mathcal{A}_{2}$), ($\mathcal{B}$) and ($\mathcal{H}_{1}$)–($\mathcal{H}_{4}$), a function $u\in W^{1,G(x,t)}(\Omega)$ is a weak solution of problem ($\mathcal{PD}$) if
$\int_{\Omega}A(x,u,Du)\cdot Dv\,\mathrm{d}x-\int_{\Omega}B(x,u,Du)v\,\mathrm{d}x=0,$ ($\mathcal{FDV}$)
for all test functions $v\in W_{0}^{1,G(x,t)}(\Omega)$.
Definition 1.8.
Under the hypotheses of Definition 1.7 and condition ($\mathcal{C}$), a function $u\in W^{1,G(x,t)}(\Omega)$ is said to be a weak solution of problem ($\mathcal{PN}$) if
$\int_{\Omega}A(x,u,Du)\cdot Dv\,\mathrm{d}x-\int_{\partial\Omega}C(x,u)v\,\mathrm{d}\sigma-\int_{\Omega}B(x,u,Du)v\,\mathrm{d}x=0,$ ($\mathcal{FNV}$)
for all $v\in W^{1,G(x,t)}(\Omega)$, where $\mathrm{d}\sigma$ denotes the surface measure on $\partial\Omega$.
Since the boundedness of weak solutions is essential for establishing their regularity, in particular, Hölder continuity, we first prove the boundedness of weak solutions to problem ($\mathcal{P}$) under both Dirichlet and Neumann conditions.
Theorem 1.9.
Assume that conditions (G0)–(G2), ($\mathcal{A}_{1}$)–($\mathcal{A}_{2}$), ($\mathcal{B}$), and ($\mathcal{H}_{1}$)–($\mathcal{H}_{4}$) hold. If $u\in W^{1,G(x,t)}(\Omega)$ is a weak solution to problem ($\mathcal{PD}$), then $u\in L^{\infty}(\Omega)$.
Theorem 1.10.
Under the assumptions of Theorem 1.9 and condition ($\mathcal{C}$), any weak solution $u\in W^{1,G(x,t)}(\Omega)$ of problem ($\mathcal{PN}$) satisfies $u\in L^{\infty}(\Omega)\cap L^{\infty}(\partial\Omega)$.
Combining the previous theorems, we establish the Hölder continuity of weak solutions to problem ($\mathcal{P}$), subject to either Dirichlet boundary conditions ($\mathcal{PD}$) or Neumann boundary conditions ($\mathcal{PN}$).
Theorem 1.11.
Assume that conditions (G0)–(G3), ($\mathcal{A}_{1}$)–($\mathcal{A}_{2}$), ($\mathcal{B}$) and ($\mathcal{H}_{1}$)–($\mathcal{H}_{4}$), hold. Let $u\in W^{1,G(x,t)}(\Omega)$ be a weak solution to the problem ($\mathcal{PD}$), and suppose that
$\max_{x\in\Omega}|u(x)|\leq M.$ ($\mathcal{M}$)
Then $u\in\mathcal{B}_{G(x,t)}(\overline{\Omega},M,\gamma,\gamma_{1},\delta)$, where
$\gamma=\gamma\big(a_{2},a_{3},a_{4},g^{-},g^{+}\big),\quad\gamma_{1}=\gamma_{1}\big(a_{1},a_{2},a_{3},a_{5},a_{6},b_{1},b_{2},b_{3},g^{+},h^{+}_{1},F_{1},M\big),\quad\delta=\min\left\{\frac{a_{4}}{4b_{3}M},2\right\}.$
Theorem 1.12.
Assume that conditions of Theorem 1.11 and condition ($\mathcal{C}$) hold. Let $u\in W^{1,G(x,t)}(\Omega)$ be a weak solution to problem ($\mathcal{PN}$), and suppose that ($\mathcal{M}$) holds. Then $u\in\mathcal{B}_{G(x,t)}(\overline{\Omega},M,\gamma,\gamma_{1},\delta)$, where
$\gamma=\gamma\big(a_{2},a_{3},a_{4},g^{-},g^{+}\big),\quad\gamma_{1}=\gamma_{1}\big(a_{1},a_{2},a_{3},a_{5},a_{6},b_{1},b_{2},b_{3},c_{1},c_{2},g^{+},h^{+}_{1},F_{1},M\big),\quad\delta=\min\left\{\frac{a_{4}}{4b_{3}M},2\right\}.$
Theorem 1.13.
Suppose that the assumptions of Theorem 1.11 or Theorem 1.12 are satisfied. Then $u\in C^{0,\alpha}(\Omega)$, where
$\alpha=\alpha\big(a_{2},a_{3},a_{4},M,n,g^{-},g^{+},F,L\big)\in(0,1].$
Theorem 1.14.
Let $u\in W^{1,G(x,t)}(\Omega)$ be a weak solution to problem ($\mathcal{PD}$) (respectively, ($\mathcal{PN}$)), under the assumptions of Theorem 1.11 (respectively, Theorem 1.12). Then $u$ is locally Hölder continuous in $\Omega$. Moreover, if the boundary $\partial\Omega$ satisfies condition (A) and the trace $u|_{\partial\Omega}$ is Hölder continuous, then $u$ is Hölder continuous up to the boundary, i.e., $u\in C^{0,\alpha}(\overline{\Omega})$ for some $\alpha\in(0,1]$.
1.3. Examples
The following discussion of customary generalized N-functions satisfying (G0)–(G3) serves to illustrate the range of problems covered by our theorems.
Example 1.1.
[$p$-growth conditions] Let $p\in(1,n)$. Consider the case where $G(x,t)=|t|^{p}$. In this setting, $G(x,t)$ is a generalized N-function that satisfies the conditions (G0)–(G3). Consequently, the main results established in this work apply to the classical $p$-growth case, which has already been extensively investigated by Ladyženskaja-Solonnikov-Uralceva [49] and G. Lieberman [50].
Example 1.2.
[$G(t)$-growth conditions] A more general framework than the classical $p$-growth is obtained by considering a generalized N-function $G(x,t)$ that does not depend on the spatial variable $x$. In this case, we write $G(t)$, and the problem reduces to the so-called Orlicz case. If
$1<g^{-}\leq\frac{g(t)t}{G(t)}\leq g^{+}<n,\ \text{for all}\ t>0,$
then, the function $G(t)$ still satisfies the assumptions (G0)–(G3), thereby falling within the scope of our results. Various aspects of this case have been previously studied in depth, notably by G. Lieberman [51].
Example 1.3.
[$p(x)$-growth conditions] Let $p:\Omega\rightarrow\mathbb{R}$ be a measurable function such that
$1<p^{-}:=\inf_{x\in\overline{\Omega}}p(x)\leq p^{+}:=\sup_{x\in\overline{\Omega}}p(x)<\infty.$ (1.6)
Assume that $p$ satisfies the log-Hölder continuity condition:
$\left|\frac{1}{p(x)}-\frac{1}{p(y)}\right|\leq\frac{c}{\ln\left(e+\frac{1}{|x-y|}\right)}\quad\text{for all }x,y\in\overline{\Omega},$ (1.7)
for some constant $c>0$. Then the function
$G(x,t)=|t|^{p(x)}$
defines a generalized N-function that satisfies the structural assumptions (G0)–(G3); see [41, Section 7.1]. Therefore, our results also apply to the variable exponent setting, which has been extensively studied by X. Fan and D. Zhao [30].
Example 1.4.
[Logarithmic growth conditions] Let $p:\Omega\rightarrow\mathbb{R}$ be a measurable function satisfying conditions (1.6)–(1.7). Then the function
$G(x,t)=|t|^{p(x)}\ln(e+|t|)$
defines a generalized N-function that satisfies the structural conditions (G0)–(G3). This example extends the case of the classical variable exponent by incorporating a logarithmic perturbation that reflects more refined material responses in heterogeneous media. This type of growth has recently attracted attention in the context of non-standard regularity theory and is encompassed by our main results.
Example 1.5.
[$(p(x),q(x))$-growth conditions] Let $p,q:\Omega\rightarrow\mathbb{R}$ be measurable functions such that
$p,q\in C^{0,1}(\overline{\Omega}),$ (1.8)
$1<p(x)<n\quad\text{and}\quad p(x)<q(x)\quad\text{for all}\ x\in\overline{\Omega},$ (1.9)
$\frac{q^{+}}{p^{-}}<1+\frac{1}{n},\quad\text{where}\quad p^{-}:=\inf_{x\in\overline{\Omega}}p(x)\ \text{and}\ q^{+}:=\sup_{x\in\overline{\Omega}}q(x),$ (1.10)
and
[…]
Under these conditions, the function
$G(x,t)=|t|^{p(x)}+|t|^{q(x)}$
defines a generalized N-function satisfying assumptions (G0)–(G3). Consequently, our results also apply to the variable exponent double phase growth setting, which has recently been the focus of active research.
Example 1.6.
[$(p,q)$-type growth conditions] Let $1<p<q<\infty$ and let $a\in L^{\infty}(\Omega)$ be a positive function satisfying
$q\leq\frac{(n+1)p}{n},$ (1.12)
as well as the Hölder continuity condition
$\frac{1}{a}\in C^{0,\frac{n}{p}(q-p)}(\Omega).$ (1.13)
In particular, the following estimate holds:
$\frac{1}{a(x)}\leq\frac{1}{a(y)}+|x-y|^{\frac{n}{p}(q-p)}\quad\text{for all }x,y\in\Omega,$ (1.14)
for some constant $c>0$. Under these conditions, the function
$G(x,t)=|t|^{p}+a(x)|t|^{q}$
defines a generalized function N satisfying the assumptions (G0)–(G3). Hence, our results also cover the semi-double phase growth case. Unfortunately, our results do not cover the classical double phase growth case, which has attracted significant attention in recent years (see  [15]).
Example 1.7.
[Logarithmic $(p,q)$-type growth conditions] Let $1<p<q<\infty$ and let $a\in L^{\infty}(\Omega)$ be a positive function satisfying the conditions (1.12)–(1.14). Then the function
$G(x,t)=|t|^{p}+a(x)|t|^{q}\ln(e+|t|)$
defines a generalized N-function that satisfies the structural conditions (G0)–(G3). This example extends the classical double-phase case by incorporating a logarithmic perturbation in the higher-order term, which reflects more refined material responses in heterogeneous media. This type of growth has recently attracted attention in the context of non-standard regularity theory and is encompassed by our main results.
Following the same structure, many additional examples of growth conditions can be generated that fall within the scope of our results.
1.4. Structure of the paper
The paper is organized as follows. In Section 2, we introduce notations and recall the fundamental definitions and properties of Musielak–Orlicz and Musielak–Orlicz–Sobolev spaces. Section 3 presents a new De Giorgi-type class and establishes key regularity results for this class, which provide the theoretical foundation for our main results, Theorems 1.2 and 1.5. Building on this framework, Section 4 proves the boundedness of weak solutions to problem ($\mathcal{P}$) in both subcritical and critical cases (Theorems 1.9 and 1.10). Furthermore, using the regularity results for the De Giorgi-type class, we establish the $C^{0,\alpha}$ regularity of weak solutions (Theorems 1.111.14).
2. Preliminaries
Here, we introduce the fundamental notations and definitions used throughout the paper. A central concept is the Musielak-Orlicz-Sobolev space, whose properties are essential for the results that follow.
2.1. Notations
The following notations will be used throughout this paper:
  • $\mathbb{R}^{n}$ denotes the $n$-dimensional Euclidean space.
  • $\Omega\subset\mathbb{R}^{n}$ is a bounded domain.
  • For a measurable set $E\subset\mathbb{R}^{n}$, $\mathrm{mes}(E)$ or $|E|$ denotes its $n$-dimensional Lebesgue measure.
  • If $u$ is a measurable function defined on $\Omega$, and $E\subset\Omega$ is measurable, we define:
    $\max_{E}u:=\mathop{\mathrm{ess\,sup}}_{x\in E}u(x),\quad\min_{E}u:=\mathop{\mathrm{ess\,inf}}_{x\in E}u(x),\quad\mathrm{osc}_{E}u:=\max_{E}u-\min_{E}u.$ (2.1)
  • For any $x_{0}\in\mathbb{R}^{n}$ and $\rho>0$, we define the open ball:
    $B_{\rho}(x_{0}):=\{x\in\mathbb{R}^{n}:|x-x_{0}|<\rho\},$
    which we may denote simply by $B_{\rho}$ when the center is clear from context. Balls $B_{\rho_{1}}$ and $B_{\rho_{2}}$ are assumed to be concentric. We denote by $\omega_{n}:=|B_{1}|$ the measure of the unit ball in $\mathbb{R}^{n}$.
  • For $0<\alpha\leq 1$, we denote:
[…]
    and
    $C^{0,\alpha}(\Omega):=\left\{u:u\in C^{0,\alpha}(\overline{\Omega^{\prime}})\text{ for all }\Omega^{\prime}\Subset\Omega\right\}.$
  • For any function $f:\Omega\rightarrow\mathbb{R}$, we define the positive part and the negative part of $f$ as:
    $f^{+}=\max\{f,0\},\quad f^{-}=\max\{-f,0\}.$
2.2. Musielak-Orlicz-Sobolev space $W^{1,G(x,t)}$
In this subsection, we recall some definitions and fundamental properties of Musielak-Orlicz and Musielak-Orlicz-Sobolev spaces. For a comprehensive bibliography on Musielak-Orlicz-Sobolev spaces, we refer the reader to [5, 24, 12, 23, 41, 28, 29, 72, 74, 75, 76, 60, 61, 62, 63, 64, 65, 66, 67, 68].
Definition 2.1.
Let $\Omega$ be an open subset of $\mathbb{R}^{n}$. A function $G:\Omega\times\mathbb{R}\to\mathbb{R}$ is called a generalized N-function if it satisfies the following conditions:
  1. (1)
    For a.a. $x\in\Omega$, the function $G(x,t)$ is even, continuous, increasing, and convex in $t$, and for each $t\in\mathbb{R}$, $G(x,t)$ is measurable in $x$;
  2. (2)
    $\lim_{t\to 0}\frac{G(x,t)}{t}=0$, for a.a. $x\in\Omega$;
  3. (3)
    $\lim_{t\to\infty}\frac{G(x,t)}{t}=\infty$, for a.a. $x\in\Omega$;
  4. (4)
    $G(x,t)>0$ for all $t>0$ and all $x\in\Omega$, and $G(x,0)=0$ for all $x\in\Omega$.
Remark 2.2 (Remark 2.1, [4]).
We give an equivalent definition of a generalized N-function that admits an integral representation. For $x\in\Omega$ and $t\geq 0$, let $g(x,t)$ be the right-hand derivative of $G(x,\cdot)$ at $t$, and define g(x,t)=g(x,t)g(x,t)=-g(x,-t) for $t<0$. Then for each $x\in\Omega$, the function $g(x,\cdot)$ is odd, real-valued, satisfies $g(x,0)=0$, and $g(x,t)>0$ for $t>0$, is right-continuous and nondecreasing on $0,+\infty)$, and
$G(x,t)=\int_{0}^{|t|}g(x,s)\,\mathrm{d}s,\quad\text{for all }x\in\Omega,\ t\in\mathbb{R}.$
Definition 2.3.
A generalized N-function $G$ satisfies the $\Delta_{2}$-condition if there exist $C_{0}>0$ and a nonnegative function $\varphi\in L^{1}(\Omega)$ such that
$G(x,2t)\leq C_{0}G(x,t)+\varphi(x),\quad\text{for a.a. }x\in\Omega\text{ and all }t\geq 0.$
Definition 2.4.
A generalized N-function $G(x,t)$ is said to satisfy:
  1. (A0)
    if there exists $\mu\in(0,1]$ such that
    $\mu\leq G^{-1}(x,1)\leq\frac{1}{\mu},\quad\text{for a.a. }x\in\mathbb{R}^{n};$
  2. (A1)
    if there exists $\mu\in(0,1]$ such that
    $\mu G^{-1}(x,t)\leq G^{-1}(y,t),$
    for every $t\in[1,1/|B|]$, for a.a. $x,y\in B$, and every ball $B\subset\mathbb{R}^{n}$ with $|B|\leq 1$;
  3. (A2)
    if for every $s>0$ there exist $\mu\in(0,1]$ and $\phi\in L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ such that
    $\mu G^{-1}(x,t)\leq G^{-1}(y,t),$
    for a.a. $x,y\in\mathbb{R}^{n}$ and for all $t\in[\phi(x)+\phi(y),s]$.
Definition 2.5.
Let $G_{1}(x,t)$ and $G_{2}(x,t)$ be two generalized N-functions.
  • (1)
    We say that $G_{1}(x,t)$ increases essentially slower than $G_{2}(x,t)$ near infinity, and we write $G_{1}\prec\prec G_{2}$, if for any $k>0$
    $\lim_{t\to\infty}\frac{G_{1}(x,kt)}{G_{2}(x,t)}=0,\quad\text{uniformly in }x\in\Omega.$
  • (2)
    We say that $G_{1}(x,t)$ is weaker than $G_{2}(x,t)$, denoted by $G_{1}\prec G_{2}$, if there exist constants $C_{1},C_{2}>0$ and a nonnegative function $\varphi\in L^{1}(\Omega)$ such that
    $G_{1}(x,t)\leq C_{1}G_{2}(x,C_{2}t)+\varphi(x),\quad\text{for a.a. }x\in\Omega\text{ and all }t\geq 0.$
Definition 2.6 (Complementary function).
For any generalized N-function $G(x,t)$, the function $\widetilde{G}:\Omega\times\mathbb{R}\to\mathbb{R}$ defined by
$\widetilde{G}(x,t):=\sup_{s\geq 0}\left(ts-G(x,s)\right),\quad\text{for all }x\in\Omega\text{ and }t\geq 0,$ (2.2)
is called the complementary function of $G$.
Remark 2.7.
From the definition of the complementary function $\widetilde{G}(x,t)$, we derive the following Young-type inequality:
$st\leq G(x,s)+\widetilde{G}(x,t),\quad\text{for all }x\in\Omega,\ s,t\geq 0.$ (2.3)
The next lemma is taken from Bahrouni–Bahrouni–Missaoui [4, Lemma 2.3].
Lemma 2.8.
Let $G(x,t)$ be a generalized N-function. Suppose that $t\mapsto g(x,t)$ is continuous and increasing on $\mathbb{R}$ for a.a. $x\in\Omega$. Moreover, assume that there exist constants $g^{-},g^{+}\in\mathbb{R}$ such that
$\widetilde{G}(x,g(x,s))$ $\leq(g^{+}-1)G(x,s),\quad\text{for all }s\geq 0,\ x\in\Omega,$ (2.4)
and
$\frac{g^{+}}{g^{+}-1}=:\widetilde{g^{-}}$ $\leq\frac{\widetilde{g}(x,s)s}{\widetilde{G}(x,s)}\leq\widetilde{g^{+}}:=\frac{g^{-}}{g^{-}-1},\quad\text{for all }x\in\Omega,\ s>0,$ (2.5)
where $\widetilde{G}(x,s)=\int_{0}^{s}\widetilde{g}(x,t)\,dt$.
Remark 2.9.
The condition (G0) implies that $G(x,t)$ and its complementary function $\widetilde{G}(x,t)$ satisfy the $\Delta_{2}$-condition.
Now, we define the Musielak-Orlicz space as follows:
$L^{G(x,t)}(\Omega):=\left\{u:\Omega\longrightarrow\mathbb{R}\ \text{measurable :}\ \rho_{G}(\lambda u)<+\infty,\ \ \text{for some}\ \lambda>0\right\},$
where
$\rho_{G}(u):=\int_{\Omega}G(x,u){\rm d}x.$ (2.6)
The space $L^{G(x,t)}(\Omega)$ is endowed with the Luxemburg norm
$\|u\|_{L^{G(x,t)}(\Omega)}:=\inf\left\{\lambda:\ \rho_{G}\left(x,\frac{u}{\lambda}\right)\leq 1\right\}.$ (2.7)
Proposition 2.10.
Let $G$ be a generalized N-function satisfies the $\Delta_{2}$-condition, then
$L^{G(x,t)}(\Omega)=\left\{u:\Omega\longrightarrow\mathbb{R}\ \text{measurable :}\ \rho_{G}(u)<+\infty\right\}.$
Proposition 2.11.
Let $G$ be a generalized N-function satisfies (G0), then the following assertions hold:
  • (1)
    $\min\{s^{g^{-}},s^{g^{+}}\}G(x,t)\leq G(x,st)\leq\max\{s^{g^{-}},s^{g^{+}}\}G(x,t),\text{ for a.a. }x\in\Omega$  $\text{ and all }s,\ t\geq 0.$
  • (2)
    $\min\{s^{\widetilde{g^{-}}},s^{\widetilde{g^{+}}}\}\widetilde{G}(x,t)\leq\widetilde{G}(x,st)\leq\max\{s^{\widetilde{g^{-}}},s^{\widetilde{g^{+}}}\}\widetilde{G}(x,t),\text{ for a.a. }x\in\Omega$  $\text{ and all }s,\ t\geq 0.$
  • (3)
    $\min\left\{\|u\|_{L^{G(x,t)}(\Omega)}^{g^{-}},\|u\|_{L^{G(x,t)}(\Omega)}^{g^{+}}\right\}\leq\rho_{G}(u)\leq\max\left\{\|u\|_{L^{G(x,t)}(\Omega)}^{g^{-}},\|u\|_{L^{G(x,t)}(\Omega)}^{g^{+}}\right\},$ for all $u\in L^{G(x,t)}(\Omega)$.
As a consequence of (2.3), we have the following result:
Lemma 2.12 (Hölder’s type inequality).
Let $\Omega$ be an open subset of $\mathbb{R}^{n}$ and $G$ be a generalized N-function satisfies (G0), then
$\left|\int_{\Omega}uv{\rm d}x\right|\leq 2\|u\|_{L^{G(x,t)}(\Omega)}\|v\|_{L^{\widetilde{G}(x,t)}(\Omega)},\ \text{for all}\ u\in L^{G(x,t)}(\Omega)\ \text{and all}\ v\in L^{\widetilde{G}(x,t)}(\Omega).$ (2.8)
The subsequent proposition deals with some topological properties of the Musielak-Orlicz space, see [72, Theorem 7.7 and Theorem 8.5].
Proposition 2.13.
  1. (1)
    Let $G(x,t)$ be a generalized N-function and $\Omega$ an open subset of $\mathbb{R}^{n}$. Then,
    1. (a)
      the space $\left(L^{G(x,t)}(\Omega),\|\cdot\|_{G}\right)$ is a Banach space;
    2. (b)
      if $G(x,t)$ satisfies (G0), then $L^{G(x,t)}(\Omega)$ is a separable and reflexive space.
  2. (2)
    Let $G_{1}(x,t)$ and $G_{2}(x,t)$ be two generalized N-functions such that $G_{1}(x,t)\prec G_{2}(x,t)$ and $\Omega$ be an open bounded subset of $\mathbb{R}^{n}$. Then, L^G_2(x,t)(Ω) ↪L^G_1(x,t)(Ω).
Now, we are ready to define the Musielak-Orlicz Sobolev space. Let $G(x,t)$ be a generalized N-function and $\Omega$ be an open subset of $\mathbb{R}^{n}$. The Musielak-Sobolev space is defined as follows
$W^{1,G(x,t)}(\Omega):=\left\{u\in L^{G(x,t)}(\Omega):\ |\nabla u|\in L^{G(x,t)}(\Omega)\right\}.$
The space $W^{1,G(x,t)}(\Omega)$ is endowed with the norm
$\|u\|:=\|u\|_{L^{G(x,t)}(\Omega)}+\|\nabla u\|_{L^{G(x,t)}(\Omega)},\ \ \text{for all}\ u\in W^{1,G(x,t)}(\Omega),$ (2.9)
where $\|\nabla u\|_{L^{G(x,t)}(\Omega)}:=\||\nabla u|\|_{L^{G(x,t)}(\Omega)}$.
We denote by $W^{1,G(x,t)}_{0}(\Omega)$ the completion of $C^{\infty}_{0}(\Omega)$ in $W^{1,G(x,t)}(\Omega)$.
Remark 2.14.
If $G$ satisfies (G0), then the space $W^{1,G(x,t)}(\Omega)$ is a reflexive and separable Banach space with respect to the norm $\|\cdot\|$.
Now we introduce the Sobolev conjugate from [6], which refines the one proposed in [13] and yields several important results presented below.
Definition 2.15.
The Sobolev conjugate of $G(x,t)$ is defined as the generalized N-function $G^{\ast}(x,t)$ given by
$G^{\ast}(x,t)=G(x,N^{-1}(x,t)),\text{ for a.a. }x\in\mathbb{R}^{n}\text{ and all }t\geq 0,$
where
$N(x,t)=\left(\int_{0}^{t}\left(\frac{\tau}{\Phi(x,\tau)}\right)^{\frac{1}{n-1}}d\tau\right)^{\frac{n-1}{n}},\text{ for }x\in\mathbb{R}^{n}\text{ and }t\geq 0.$
Proposition 2.16 ([6]).
Let $G(x,t)$ be a generalized N-function that satisfies (G0)–(G2) and $G^{*}(x,t)$ its Sobolev conjugate function, then the following assertions hold:
  • (1)
    $\min\left\{s^{g^{-}_{*}},s^{g^{+}_{*}}\right\}G^{*}(x,t)\leq G^{*}(x,st)\leq\max\left\{s^{g^{-}_{*}},s^{g^{+}_{*}}\right\}G^{*}(x,t),\text{ for a.a. }x\in\Omega$  $\text{ and all }s,\ t\geq 0$;
  • (2)
    ${1<g^{-}_{*}\leq\frac{g^{*}(x,t)t}{G^{*}(x,t)}\leq g^{+}_{*},\ \ \text{for all}\ x\in\Omega\ \text{and all}\ t>0}$;
  • (3)
    $\min\left\{\|u\|_{L^{G^{*}(x,t)}(\Omega)}^{g^{-}_{*}},\|u\|_{L^{G^{*}(x,t)}(\Omega)}^{g^{+}_{*}}\right\}\leq\rho_{G^{*}}(u)\leq\max\left\{\|u\|_{L^{G^{*}(x,t)}(\Omega)}^{g^{-}_{*}},\|u\|_{L^{G^{*}(x,t)}(\Omega)}^{g^{+}_{*}}\right\},$ for all $u\in L^{G^{*}(x,t)}(\Omega)$;
where ${g_{*}^{-}:=\frac{ng^{-}}{n-g^{-}}}$, ${g_{*}^{+}:=\frac{ng^{+}}{n-g^{+}}}$ and ${G^{*}(x,t)=\int_{0}^{t}g^{*}(x,s){\rm d}s}$.
Proposition 2.17 ([14]).
Let $\Omega$ be a bounded domain in $\mathbb{R}^{n}$ and $G(x,t)$ be a generalized N-function that satisfies (G1)–(G2). Let $\vartheta(x,t)$ be a generalized N-function such that
$\vartheta(x,t)\prec\prec G^{*}(x,t),\text{ and }\int_{\Omega}\vartheta(x,t)\,\mathrm{d}x<\infty,\ \text{for }t>0.$ (2.10)
  1. (1)
    The embedding
    $W^{1,G(x,t)}_{0}(\Omega)\hookrightarrow L^{\vartheta(x,t)}(\Omega)$ (2.11)
    is compact.
  2. (2)
    Assume, in addition, that $\Omega$ is a bounded domain with a Lipschitz boundary $\partial\Omega$. Then, the embedding
    $W^{1,G(x,t)}(\Omega)\hookrightarrow L^{\vartheta(x,t)}(\Omega)$ (2.12)
    is compact.
Remark 2.18.
If the function $G(x,t)$ satisfies assumptions (G0)–(G2), then it also fulfills conditions (A0)–(A1) as stated in L. Diening and A. Cianchi paper [14], and it satisfies condition (A2) due to the boundedness of $\Omega\subset\mathbb{R}^{n}$. Consequently, the embedding result given in Proposition 2.17 remains valid under assumptions (G0)–(G2).
3. A New Class of De Giorgi-Type Functions $\mathcal{B}_{G(x,t)}$
In this section, we introduce and analyze a new De Giorgi-type class of functions denoted by $\mathcal{B}_{G(x,t)}$. Our main results are encapsulated in Theorems 1.2 and 1.5, which establish key regularity properties for functions belonging to this class.
The proof strategies we adopt are inspired by the foundational works of X. Fan [30] and G. Lieberman [51]. However, our approach involves significant technical refinements that are essential to overcome the challenges arising from the nonstandard growth behavior and spatial heterogeneity inherent in the Musielak–Orlicz framework. In particular, the interaction between Orlicz-type growth and variable exponent dependence introduces substantial analytical difficulties, which require new ideas and careful adaptations of classical methods.
These extensions allow us to unify and generalize existing results under a set of natural and flexible assumptions, thereby providing a broader and more robust regularity theory applicable to a wide range of problems.
We begin by collecting several auxiliary lemmas that will play a crucial role in the subsequent analysis.
Lemma 3.1 (see Lemma 4.7 of Chap. II of [49]).
Let $\{Y_{h}\}_{h\geq 0}$ be a sequence of non-negative numbers satisfying the recursion relation
$Y_{h+1}\leq cb^{h}Y_{h}^{1+\varepsilon},\quad h=0,1,2,\dots,$
where $c>0$, $b>1$, and $\varepsilon>0$. If
$Y_{0}\leq\theta:=c^{-\frac{1}{\varepsilon}}b^{-\frac{1}{\varepsilon^{2}}},$
then
$Y_{h}\leq\theta b^{-\frac{h}{\varepsilon}},\quad\forall h=0,1,2\ldots,$
and hence $Y_{h}\to 0$ as $h\to\infty$.
Lemma 3.2 (see Lemma 3.5 of Chap. II of [49]).
For any $u\in W^{1,1}(B_{\rho})$ and arbitrary numbers $k$ and $l$ with $l>k$, the following inequality holds:
$(l-k)|A_{l,\rho}|^{1-\frac{1}{n}}\leq\beta\frac{\rho^{n}}{|B_{\rho}\setminus A_{k,\rho}|}\int_{A_{k,\rho}\setminus A_{l,\rho}}|Du|\,{\rm d}x,$ (3.1)
where $A_{k,\rho}=\{x\in B_{\rho}:u(x)>k\}$, $\beta=\beta(n)>1$ is a constant depending only on $n$.
Lemma 3.3 (see Lemma 2.3 of [30] or Lemma 4.8 of Chap. II of [49]).
Suppose a function $u(x)$ is measurable and bounded in some ball $B_{R_{0}}$. Consider balls $B_{R}$ and $B_{bR}$ which have a common center with $B_{R_{0}}$, where $b>1$ is a fixed constant, and suppose that for any $0<R\leq b^{-1}R_{0}$, at least one of the following two inequalities is valid:
$\operatorname{osc}_{B_{R}}u\leq c_{1}R^{\varepsilon},\quad\text{or}\quad\operatorname{osc}_{B_{R}}u\leq\theta\operatorname{osc}_{B_{bR}}u,$
where $c_{1},\varepsilon\leq 1$ and $\theta<1$ are positive constants. Then $u\in C^{0,\alpha}(B_{R_{0}})$, where $\alpha=\min\{\varepsilon,-\log_{b}\theta\}$.
The following lemmas are fundamental to our subsequent analysis.
Lemma 3.4.
Under assumptions (G0)–(G1) and (G3), for any $x_{0}\in\Omega$, $t>1$, and $R\in(0,1)$, define
$f^{-}(R)$ $:=\inf_{x\in B_{R}(x_{0})}r(x,t),$
$f^{+}(R)$ $:=\sup_{x\in B_{R}(x_{0})}r(x,t).$
Then the following properties hold.
  1. $(1)$
    For $t>1$, osc_B_Rr(x,t)=f^+(R)-f^-(R)=η(R).
  2. $(2)$
    For all $x\in B_{R}(x_{0})$ and $t>1$,
    $g^{-}\leq f^{-}(R)\leq r(x,t)\leq f^{+}(R)\leq g^{+}.$
  3. $(3)$
    For all $x\in B_{R}(x_{0})$ and $t>1$,
    $F^{-1}t^{f^{-}(R)}\leq G(x,t)\leq Ft^{f^{+}(R)}.$
  4. $(4)$
    For all sufficiently small $R>0$,
    $R^{f^{-}(R)-f^{+}(R)}\leq L,\quad\text{where }L:=\exp(L_{0}).$
  5. $(5)$
    The oscillation decays to zero:
    $\lim_{R\to 0^{+}}\left(f^{+}(R)-f^{-}(R)\right)=0.$
Proof.
Proof of (1): Let $t>1$. We first show that
$\eta(R)\leq f^{+}(R)-f^{-}(R).$
For any $x,y\in B_{R}(x_{0})$, we have
$r(x,t)\leq f^{+}(R)\quad\text{and}\quad r(y,t)\geq f^{-}(R),$
so that
$r(x,t)-r(y,t)\leq f^{+}(R)-f^{-}(R).$
Similarly,
$r(y,t)-r(x,t)\leq f^{+}(R)-f^{-}(R),$
and hence
$|r(x,t)-r(y,t)|\leq f^{+}(R)-f^{-}(R).$
Taking the supremum over all $x,y\in B_{R}(x_{0})$, it yields that
$\eta(R)=\sup_{x,y\in B_{R}(x_{0})}|r(x,t)-r(y,t)|\leq f^{+}(R)-f^{-}(R).$
Next, we show that
$f^{+}(R)-f^{-}(R)\leq\eta(R).$
Let $\varepsilon>0$ be arbitrary. By the definitions of supremum and infimum, there exist points $x_{\varepsilon},y_{\varepsilon}\in B_{R}(x_{0})$ such that
$r(x_{\varepsilon},t)>f^{+}(R)-\frac{\varepsilon}{2}\quad\text{and}\quad r(y_{\varepsilon},t)<f^{-}(R)+\frac{\varepsilon}{2}.$
Then,
$r(x_{\varepsilon},t)-r(y_{\varepsilon},t)>\left(f^{+}(R)-\frac{\varepsilon}{2}\right)-\left(f^{-}(R)+\frac{\varepsilon}{2}\right)=f^{+}(R)-f^{-}(R)-\varepsilon.$
Since $|r(x_{\varepsilon},t)-r(y_{\varepsilon},t)|\geq r(x_{\varepsilon},t)-r(y_{\varepsilon},t)$, we have
$|r(x_{\varepsilon},t)-r(y_{\varepsilon},t)|>f^{+}(R)-f^{-}(R)-\varepsilon.$
But by the definition of $\eta(R)$,
$\eta(R)\geq|r(x_{\varepsilon},t)-r(y_{\varepsilon},t)|>f^{+}(R)-f^{-}(R)-\varepsilon.$
Since $\varepsilon>0$ is arbitrary, it follows that
$\eta(R)\geq f^{+}(R)-f^{-}(R).$
Combining the two inequalities, we obtain
$\eta(R)=f^{+}(R)-f^{-}(R).$
By definition, $\operatorname{osc}_{B_{R}}r(x,t)=f^{+}(R)-f^{-}(R)$, so
$\operatorname{osc}_{B_{R}}r(x,t)=f^{+}(R)-f^{-}(R)=\eta(R),$
as desired.
Proof of (2): From the definitions of the infimum and supremum, we immediately obtain
$f^{-}(R)\leq r(x,t)\leq f^{+}(R),\qquad\forall\,x\in B_{R}(x_{0}).$
Since $B_{R}(x_{0})\subset\omega$, this establishes the desired chain of inequalities.
Proof of (3): Fix $x\in B_{R}(x_{0})$ and $t>1$. Consider the function $h(s)=\ln G(x,s)$. Its derivative is
$h^{\prime}(s)=\frac{g(x,s)}{G(x,s)}=\frac{r(x,s)}{s}.$
From part (2), we have $f^{-}(R)\leq r(x,s)\leq f^{+}(R)$ for all $s>1$. This implies the differential inequalities
$\frac{f^{-}(R)}{s}\leq h^{\prime}(s)\leq\frac{f^{+}(R)}{s}.$
Integrating these inequalities from $s=1$ to $s=t$, it yields ttha
$f^{-}(R)\ln t$ $\leq\ln G(x,t)-\ln G(x,1)\leq f^{+}(R)\ln t.$
Exponentiating and using the bounds $F^{-1}\leq G(x,1)\leq F$ from (G1), gives the desired result
$F^{-1}t^{f^{-}(R)}\leq G(x,t)\leq Ft^{f^{+}(R)}.$
Proof of (4): From (G3) and $(1)$, we have
$\eta(R)\leq\frac{L_{0}}{|\ln(2R)|},$
hence, for $R<\frac{1}{2}$, since $|\ln(2R)|>|\ln R|$, one has
$\eta(R)<\frac{L_{0}}{|\ln R|}.$ (3.2)
Consequently,
$R^{f^{-}(R)-f^{+}(R)}$ $=\exp\left(-\eta(R)\ln R\right)=\exp\left(\eta(R)|\ln R|\right)$
$\leq\exp\left(\frac{L_{0}}{|\ln R|}\cdot|\ln R|\right)=\exp(L_{0})=L.$
Proof of (5): From (G3), it follows directly that
$0\leq\lim_{R\to 0^{+}}\eta(R)\leq\lim_{R\to 0^{+}}\frac{L_{0}}{|\ln(2R)|}=0,$
which completes the proof. ∎
Lemma 3.5.
Assume (G0)–(G1) and (G3) hold. For any $M\geq 1$, there exists $R_{M}>0$ such that for all $R<R_{M}$,
$M^{f^{+}(R)-f^{-}(R)}\leq 2.$
Specifically, one may take $R_{M}=\min\left\{R_{0},\,\exp\left(-\frac{L_{0}\ln M}{\ln 2}\right)\right\}$, where $R_{0}$ and $L_{0}$ are the constants from (G3).
Proof.
Recall, from Lemma 3.4, that $\eta(R)=f^{+}(R)-f^{-}(R)$. By (3.2), we have $\eta(R)\leq\frac{L_{0}}{|\ln R|}$ for $R<R_{0}$. Consequently,
$M^{\eta(R)}\leq\exp\left(\frac{L_{0}\ln M}{|\ln R|}\right).$
Choose $R_{M}=\min\left\{R_{0},\,\exp\left(-\frac{L_{0}\ln M}{\ln 2}\right)\right\}$. Then, for all $R<R_{M}$,
$\frac{L_{0}\ln M}{|\ln R|}\leq\ln 2,$
which implies
$M^{\eta(R)}\leq\exp(\ln 2)=2.$
Remark 3.6.
In view of assumption (G3), together with Lemma 3.3 and Lemma 3.4-(1), the regularity of weak solutions to Problem ($\mathcal{P}$) is strongly influenced by the behavior of the function $r(x,t)$, in particular by its oscillation and regularity properties. This highlights the central role of assumption (G3).
Proof of Theorem 1.2.
Let $u\in\mathcal{B}_{G(x,t)}(\Omega,M,\gamma,\gamma_{1},\delta)$, where $G(x,t)$ satisfies conditions (G0)–(G1) and (G3). Without loss of generality, assume $M\geq 1$, $\gamma\geq 1$ and $L\geq 1$.
To prove Theorem 1.2, it suffices to show that for every point $x_{0}\in\Omega$, there exists a ball $B_{R_{0}}(x_{0})\subset\Omega$ such that $u\in C^{0,\alpha}(B_{R_{0}}(x_{0}))$, where $\alpha$ is the constant stated in the theorem.
Fix $x_{0}\in\Omega$ and choose $R_{0}\in(0,R_{M})$ where $R_{M}$ is from Lemma 3.5, such that $\overline{B}_{R_{0}}(x_{0})\subset\Omega$ and $R_{M}<1$. Then, for an arbitrary $R\in(0,R_{0}]$, at least one of the functions $w:=\pm u$ satisfies the following condition.
$\left|\left\{x\in B_{\frac{R}{2}}:w(x)>\max_{B_{R}}w(x)-\tfrac{1}{2}\mathrm{osc}_{B_{R}}u\right\}\right|\leq\tfrac{1}{2}\,|B_{\frac{R}{2}}|.$ (3.3)
From this point on, we denote by $w$ the function (either $u$ or u-u) that satisfies inequality (3.3). Define
$\tau:=\max\left\{2,\frac{2}{\delta}\right\},\quad\psi:=\tau^{-1}\,\mathrm{osc}_{B_{R}}u,$ (3.4)
and
$k^{\prime}:=\max_{B_{R}}w(x)-\psi.$ (3.5)
It follows that
$k^{\prime}\geq\max_{B_{R}}w(x)-\tfrac{1}{2}\,\mathrm{osc}_{B_{R}}u,$ (3.6)
and
$k^{\prime}\geq\max_{B_{R}}w(x)-\delta M.$ (3.7)
Remark 3.7.
Note that inequality (3.7) guarantees that conditions (1.2) and (1.3) hold for all $k\geq k^{\prime}$ and for all $\rho\leq R$.
Under these assumptions and notations, the proof of Theorem 1.2 is completed through the lemmas presented below, specifically Lemmas 3.8 to 3.11.
Lemma 3.8.
There is a positive constant $\theta=\theta(n,g^{+},L,F,\gamma)$ such that if
$|A_{k^{0},\frac{R}{2}}|\leq\theta R^{n},$ (3.8)
then at least one of the following two inequalities holds:
$H\leq\left(\frac{\gamma+\gamma_{1}+1}{\gamma}\right)^{\frac{1}{f^{+}(R)}}R,$ (3.9)
or
$\max_{B_{\frac{R}{4}}}w(x)\leq k^{0}+\frac{H}{2},$ (3.10)
where
$0<H<\psi,\quad k^{0}=\max_{B_{R}}w(x)-H.$
Proof.
If inequality (3.9) does not hold, i.e., if we assume
$\left(\frac{\gamma+\gamma_{1}+1}{\gamma}\right)<\left(\frac{H}{R}\right)^{f^{+}(R)}.$ (3.11)
For $h=0,1,2,\ldots$, set
$\rho_{h}=\frac{R}{4}+\frac{R}{2^{h+2}},\quad k_{h}=k^{0}+\frac{H}{2}-\frac{H}{2^{h+1}},$
$Y_{h}=R^{-n}|A_{k_{h},\rho_{h}}|,\quad D_{h+1}=A_{k_{h},\rho_{h+1}}\backslash A_{k_{h+1},\rho_{h+1}}.$
Claim : $Y_{h}\to 0$ as $h\to\infty$.
Applying inequality (1.2) to $k=k_{h}$, $\rho=\rho_{h}$ and $\rho-\sigma\rho=\rho_{h+1}$ for $h=0,1,2,\ldots$, and using Proposition 2.11 and Lemma 3.4 (with $t=\frac{H}{R}>1$), we obtain
$\begin{split}\int_{A_{k_{h},\rho_{h+1}}}G(x,|Dw|){\rm d}x{\rm d}x&\leq\gamma\int_{A_{k_{h},\rho_{h}}}G\left(x,\left|\frac{w(x)-k_{h}}{\rho_{h}-\rho_{h+1}}\right|\right){\rm d}x+\gamma_{1}|A_{k_{h},\rho_{h}}|\\ &\leq\gamma\int_{A_{k_{h},\rho_{h}}}G\left(x,\left|\frac{\max_{B_{R}}w(x)-\left(k^{0}+\frac{H}{2}-\frac{H}{2^{h+1}}\right)}{\frac{R}{2^{h+2}}-\frac{R}{2^{h+3}}}\right|\right){\rm d}x+\gamma_{1}|A_{k_{h},\rho_{h}}|\\ &\leq\gamma\int_{A_{k_{h},\rho_{h}}}G\left(x,\frac{\frac{H}{2}+\frac{H}{2^{h+1}}}{\frac{R}{2^{h+3}}}\right){\rm d}x+\gamma_{1}|A_{k_{h},\rho_{h}}|\\ &\leq\gamma\int_{A_{k_{h},\rho_{h}}}G\left(x,\frac{\frac{(2^{h}+1)H}{2^{h+1}}}{\frac{R}{2^{h+3}}}\right){\rm d}x+\gamma_{1}|A_{k_{h},\rho_{h}}|\\ &\leq\gamma\int_{A_{k_{h},\rho_{h}}}G\left(x,\frac{2^{h+3}H}{R}\right){\rm d}x+\gamma_{1}|A_{k_{h},\rho_{h}}|\\ &\leq\gamma\left(2^{h+3}\right)^{g^{+}}\int_{A_{k_{h},\rho_{h}}}G\left(x,\frac{H}{R}\right){\rm d}x+\gamma_{1}|A_{k_{h},\rho_{h}}|\\ &\leq\gamma\left[\left(2^{h+3}\right)^{g^{+}}\int_{A_{k_{h},\rho_{h}}}F\left(\frac{H}{R}\right)^{f^{+}(R)}{\rm d}x+\frac{\gamma_{1}}{\gamma}|A_{k_{h},\rho_{h}}|\right]\\ &\leq\gamma F\left[\left(2^{h+3}\right)^{g^{+}}H^{f^{+}(R)}R^{-f^{+}(R)}|A_{k_{h},\rho_{h}}|+\frac{\gamma_{1}}{\gamma}|A_{k_{h},\rho_{h}}|\right].\end{split}$ (3.12)
In light of Lemma 3.4, we see that
$\begin{split}\int_{A_{k_{h},\rho_{h+1}}}G(x,|Dw|){\rm d}x&\geq F^{-1}\left[\int_{A_{k_{h},\rho_{h+1}}}|Dw|^{f^{-}(R)}{\rm d}x-|A_{k_{h},\rho_{h+1}}|\right]\\ &\geq F^{-1}\left[\int_{D_{h+1}}|Dw|^{f^{-}(R)}{\rm d}x-|A_{k_{h},\rho_{h}}|\right].\end{split}$ (3.13)
Therefore, by inequality (3.12), it follows that
$\begin{split}\int_{D_{h+1}}|Dw|^{f^{-}(R)}{\rm d}x&\leq\gamma F^{2}\left[\left(2^{h+3}\right)^{g^{+}}H^{f^{+}(R)}R^{-f^{+}(R)}|A_{k_{h},\rho_{h}}|+\frac{\gamma_{1}}{\gamma}|A_{k_{h},\rho_{h}}|\right]+|A_{k_{h},\rho_{h}}|\\ &\leq\gamma F^{2}\left[\left(2^{h+3}\right)^{g^{+}}H^{f^{+}(R)}R^{-f^{+}(R)}|A_{k_{h},\rho_{h}}|+\frac{\gamma+\gamma_{1}+1}{\gamma}|A_{k_{h},\rho_{h}}|\right]\\ &\leq\gamma F^{2}\left[\left(2^{h+3}\right)^{g^{+}}H^{f^{+}(R)}R^{-f^{+}(R)}+\frac{\gamma+\gamma_{1}+1}{\gamma}\right]|A_{k_{h},\rho_{h}}|\\ &\leq\gamma F^{2}\left[\left(2^{h+3}\right)^{g^{+}}H^{f^{+}(R)}R^{-f^{+}(R)}+\frac{\gamma+\gamma_{1}+1}{\gamma}\right]R^{n}Y_{h}.\end{split}$ (3.14)
Hence, by inequalities (3.11) and (3.14), we have
$\begin{split}\int_{D_{h+1}}|Dw|^{f^{-}(R)}{\rm d}x&\leq\gamma F^{2}\left[\left(2^{h+3}\right)^{g^{+}}H^{f^{+}(R)}R^{-f^{+}(R)}+\frac{\gamma+\gamma_{1}+1}{\gamma}\right]R^{n}Y_{h}\\ &\leq\gamma F^{2}\left[\left(2^{h+3}\right)^{g^{+}}H^{f^{+}(R)}R^{-f^{+}(R)}+\left(HR^{-1}\right)^{f^{+}(R)}\right]R^{n}Y_{h}\\ &\leq\gamma F^{2}\left(2^{h+4}\right)^{g^{+}}H^{f^{+}(R)}R^{n-f^{+}(R)}Y_{h}.\end{split}$ (3.15)
Consequently, using Hölder’s inequality and inequality (3.15), we obtain
$\begin{split}\int_{D_{h+1}}|Dw|{\rm d}x&\leq\left(\int_{D_{h+1}}|Dw|^{f^{-}(R)}{\rm d}x\right)^{\frac{1}{f^{-}(R)}}|D_{h+1}|^{\frac{f^{-}(R)-1}{f^{-}(R)}}\\ &\leq\left(\gamma F^{2}\left(2^{h+4}\right)^{g^{+}}H^{f^{+}(R)}R^{n-f^{+}(R)}Y_{h}\right)^{\frac{1}{f^{-}(R)}}(R^{n}Y_{h})^{\frac{f^{-}(R)-1}{f^{-}(R)}}\\ &=\left(\gamma F^{2}\left(2^{h+4}\right)^{g^{+}}H^{f^{+}(R)}\right)^{\frac{1}{f^{-}(R)}}R^{\frac{nf^{-}(R)-f^{+}(R)}{f^{-}(R)}}Y_{h}.\end{split}$ (3.16)
Applying inequality (3.1) to $k=k_{h}$, $l=k_{h+1}$ and $\rho=\rho_{h+1}$, we get
$\begin{split}\int_{D_{h+1}}|Dw|{\rm d}x&\geq\beta^{-1}(k_{h+1}-k_{h})(R^{n}Y_{h+1})^{1-\frac{1}{n}}\rho_{h+1}^{-n}|B_{\rho_{h+1}}\backslash A_{k_{h},\rho_{h+1}}|\\ &\geq\beta^{-1}2^{-(h+2)}HR^{n-1}Y_{h+1}^{1-\frac{1}{n}}2^{n}R^{-n}\left(|B_{\frac{R}{4}}|-|A_{k^{0},\frac{R}{2}}|\right).\end{split}$ (3.17)
If inequality (3.8) holds and
$\theta\leq\frac{1}{2}\cdot 4^{-n}\omega_{n},$
then
$|B_{R/4}|-|A_{k^{0},\frac{R}{2}}|\geq 4^{-n}R^{n}\omega_{n}-\frac{1}{2}\cdot 4^{-n}\omega_{n}R^{n}=2^{-2n-1}\omega_{n}R^{n}$
and consequently from inequality (3.17), it follows that
$\int_{D_{h+1}}|Dw|{\rm d}x\geq\beta^{-1}2^{-(h+3+n)}HR^{n-1}\omega_{n}Y_{h+1}^{1-\frac{1}{n}}.$ (3.18)
From inequalities (3.16), (3.18) and the fact that
$0<H<\psi\leq M,\ M\geq 1,\ \gamma\geq 1,\ L\geq 1,\ R<1,$
we deduce, by Lemma 3.5, that
$\begin{split}Y_{h+1}^{1-\frac{1}{n}}&\leq\beta 2^{h+3+n}H^{-1}R^{1-n}\omega_{n}^{-1}\int_{D_{h+1}}|Dw|{\rm d}x\\ &\leq\beta 2^{h+3+n}H^{-1}R^{1-n}\omega_{n}^{-1}\left(\gamma F^{2}\left(2^{h+4}\right)^{g^{+}}H^{f^{+}(R)}\right)^{\frac{1}{f^{-}(R)}}R^{\frac{nf^{-}(R)-f^{+}(R)}{f^{-}(R)}}Y_{h}\\ &=\beta 2^{h+3+n}\omega_{n}^{-1}H^{\frac{f^{+}(R)-f^{-}(R)}{f^{-}(R)}}\left(\gamma F^{2}\left(2^{h+4}\right)^{g^{+}}\right)^{\frac{1}{f^{-}(R)}}R^{\frac{nf^{-}(R)-f^{+}(R)}{f^{-}(R)}-n+1}Y_{h}\\ &\leq\beta 2^{h+3+n}\omega_{n}^{-1}M^{\frac{f^{+}(R)-f^{-}(R)}{f^{-}(R)}}\left(\gamma F^{2}\left(2^{h+4}\right)^{g^{+}}\right)^{\frac{1}{f^{-}(R)}}R^{\frac{f^{-}(R)-f^{+}(R)}{f^{-}(R)}}Y_{h}\\ &\leq\gamma F^{2}\left(2^{h+4}\right)^{g^{+}}\beta 2^{h+3+n}\omega_{n}^{-1}2^{\frac{1}{f^{-}(R)}}R^{f^{-}(R)-f^{+}(R)}Y_{h}\\ &\leq\gamma F^{2}\left(2^{h+4}\right)^{g^{+}}\beta 2^{h+4+n}\omega_{n}^{-1}R^{f^{-}(R)-f^{+}(R)}Y_{h}\\ &\leq 2^{(g^{+}+1)h}\gamma F^{2}\beta 2^{4g^{+}+4+n}\omega_{n}^{-1}R^{f^{-}(R)-f^{+}(R)}Y_{h}\\ &\leq 2^{(g^{+}+1)h}\gamma F^{2}\beta 2^{4g^{+}+4+n}\omega_{n}^{-1}LY_{h}\\ &\leq c_{1}b_{1}^{h}Y_{h},\end{split}$ (3.19)
where $b_{1}=2^{(g^{+}+1)}>1$, $c_{1}=\gamma F^{2}\beta 2^{4g^{+}+4+n}\omega_{n}^{-1}L+1>1$ and $\beta=\beta(n)$ as in Lemma 3.2. Hence
$Y_{h+1}\leq cb^{h}Y_{h}^{1+\frac{1}{n-1}}=cb^{h}Y_{h}^{1+\varepsilon},$
where $b=b_{1}^{\frac{n}{n-1}}=b(n,g^{+})>1$, $c=c_{1}^{\frac{n}{n-1}}=c(n,g^{+},L,F,\gamma)$ and $\varepsilon=\frac{1}{n-1}$. Now we choose
$\theta=\min\left\{\frac{1}{2}\cdot 4^{-n}\omega_{n},\quad c^{-\frac{1}{\varepsilon}}b^{-\frac{1}{\varepsilon^{2}}}\right\}.$
Then $\theta=\theta(n,g^{+},L,F,\gamma)$ and when
$|A_{k^{0},\frac{R}{2}}|\leq\theta R^{n}\quad\text{and}\quad H>\left(\frac{\gamma+\gamma_{1}+1}{\gamma}\right)^{\frac{1}{f^{+}(R)}}R,$
we have
$Y_{0}=R^{-n}|A_{k^{0},\frac{R}{2}}|\leq\theta\leq c^{-\frac{1}{\varepsilon}}b^{-\frac{1}{\varepsilon^{2}}}$
and consequently, by Lemma 3.1, $Y_{h}\to 0$ as $h\to\infty$. This shows the Claim.
From the claim, it yields that
$\max_{B_{\frac{R}{4}}}w(x)\leq\lim_{h\to\infty}k_{h}=k^{0}+\frac{H}{2}.$
Lemma 3.8 is proved. ∎
Lemma 3.9.
For any given $\theta>0$, there is a natural number $s=s(n,g^{-},g^{+},F,L,\beta,\gamma)>2$ such that either
$\psi\leq 2^{s}\left(\frac{\gamma+\gamma_{1}+1}{\gamma}\right)^{\frac{1}{f^{+}(R)}}R$ (3.20)
or
$|A_{k^{0},\frac{R}{2}}|\leq\theta R^{n}$ (3.21)
holds, where
$k^{0}=\max_{B_{R}}w(x)-2^{-s+1}\psi.$
Proof.
If inequality (3.20) does not hold, i.e., if we assume
$\psi>2^{s}\left(\frac{\gamma+\gamma_{1}+1}{\gamma}\right)^{\frac{1}{f^{+}(R)}}R,$ (3.22)
where the natural number $s$ will be chosen later, we proceed as follows.
For $t=0,\ldots,s-1$, define
$k_{t}=\max_{B_{R}}w(x)-2^{-t}\psi,\quad D_{t}=A_{k_{t},\frac{R}{2}}\setminus A_{k_{t+1},\frac{R}{2}}.$
By (3.22), one has
$\frac{2^{-t+1}\psi}{R}>1,\quad\text{for }t=0,\ldots,s-1.$ (3.23)
Applying inequality (1.2) to $\rho=R$, $\rho-\sigma\rho=\frac{R}{2}$ and $k=k_{t}$ for $t=0,\ldots,s-2$, and using Proposition 2.11, Lemma  3.4 (with $t=\frac{2^{-t+1}\psi}{R}$), we obtain
$\begin{split}\int_{A_{k_{t},\frac{R}{2}}}G(x,|Dw|){\rm d}x&\leq\gamma\left[\int_{A_{k_{t},R}}G\left(x,\frac{\omega(x)-k_{t}}{\frac{R}{2}}\right){\rm d}x+\frac{\gamma_{1}}{\gamma}|A_{k_{t},R}|\right]\\ &\leq\gamma\left[\int_{A_{k_{t},R}}G\left(x,\frac{2^{-t}\psi}{\frac{R}{2}}\right){\rm d}x+\frac{\gamma_{1}}{\gamma}|A_{k_{t},R}|\right]\\ &\leq\gamma\left[\int_{A_{k_{t},R}}G\left(x,\frac{2^{-t+1}\psi}{R}\right){\rm d}x+\frac{\gamma_{1}}{\gamma}|A_{k_{t},R}|\right]\\ &\leq\gamma\left[\int_{A_{k_{t},R}}F\left(\frac{2^{-t+1}\psi}{R}\right)^{f^{+}(R)}{\rm d}x+\frac{\gamma_{1}}{\gamma}|A_{k_{t},R}|\right]\\ &\leq\gamma F\left[\left(2^{-t+1}\psi\right)^{f^{+}(R)}R^{-f^{+}(R)}|A_{k_{t},R}|+\frac{\gamma_{1}}{\gamma}|A_{k_{t},R}|\right].\end{split}$ (3.24)
On the other hand, from Lemma 3.4, we deduce that
$\begin{split}\int_{A_{k_{t},\frac{R}{2}}}G(x,|Dw|){\rm d}x&\geq F^{-1}\left[\int_{A_{k_{t},\frac{R}{2}}}|Dw|^{f^{-}(R)}{\rm d}x-|A_{k_{t},\frac{R}{2}}|\right]\\ &\geq F^{-1}\left[\int_{D_{t}}|Dw|^{f^{-}(R)}{\rm d}x-|A_{k_{t},R}|\right].\end{split}$ (3.25)
Thus, combining (3.22), (3.24), and (3.25), we get
$\begin{split}\int_{D_{t}}|Dw|^{f^{-}(R)}{\rm d}x&\leq\gamma F^{2}\left[\left(2^{-t+1}\psi\right)^{f^{+}(R)}R^{-f^{+}(R)}|A_{k_{t},R}|+\frac{\gamma_{1}}{\gamma}|A_{k_{t},R}|\right]+|A_{k_{t},R}|\\ &\leq\gamma F^{2}\left[\left(2^{-t+1}\psi\right)^{f^{+}(R)}R^{-f^{+}(R)}|A_{k_{t},R}|+\frac{\gamma+\gamma_{1}+1}{\gamma}|A_{k_{t},R}|\right]\\ &\leq\gamma F^{2}\left[\left(2^{-t+1}\psi\right)^{f^{+}(R)}R^{-f^{+}(R)}+\left(2^{-s}\psi R^{-1}\right)^{f^{+}(R)}\right]|A_{k_{t},R}|\\ &\leq\gamma F^{2}\left[\left(2^{-t+1}\psi\right)^{f^{+}(R)}R^{-f^{+}(R)}+\left(2^{-s}\psi R^{-1}\right)^{f^{+}(R)}\right]\omega_{n}R^{n}\\ &\leq\gamma F^{2}\left[\left(2^{-t+1}\psi\right)^{f^{+}(R)}+\left(2^{-t}\psi\right)^{f^{+}(R)}\right]\omega_{n}R^{n-f^{+}(R)}\\ &\leq\gamma F^{2}2^{f^{+}(R)+1}\omega_{n}\left(2^{-t}\psi\right)^{f^{+}(R)}R^{n-f^{+}(R)}\\ &\leq\gamma F^{2}2^{g^{+}+1}\omega_{n}\left(2^{-t}\psi\right)^{f^{+}(R)}R^{n-f^{+}(R)},\quad(t=0,\ldots,s-2).\end{split}$ (3.26)
Now, applying inequality (3.1) to $w$ with $k=k_{t}$, $l=k_{t+1}$, $\rho=\frac{R}{2}$, for $t=0,\ldots,s-2$, we obtain
[…]
It follows, from inequalities (3.26), (3.27) and Lemma 3.5, that
$\begin{split}|A_{k_{s-1},\frac{R}{2}}|^{\frac{(n-1)f^{-}(R)}{n}}&\leq(4\beta\omega_{n}^{-1}2^{t}\psi^{-1})^{f^{-}(R)}\int_{D_{t}}|Dw|^{f^{-}(R)}{\rm d}x\cdot|D_{t}|^{f^{-}(R)-1}\\ &\leq(4\beta\omega_{n}^{-1}2^{t}\psi^{-1})^{f^{-}(R)}\gamma F^{2}2^{g^{+}+1}\omega_{n}\left(2^{-t}\psi\right)^{f^{+}(R)}R^{n-f^{+}(R)}|D_{t}|^{f^{-}(R)-1}\\ &\leq(2^{n}\beta\omega_{n}^{-1})^{g^{+}}\gamma F^{2}2^{g^{+}+1}\omega_{n}\left(2^{-t}\psi\right)^{f^{+}(R)-f^{-}(R)}R^{n-f^{+}(R)}|D_{t}|^{f^{-}(R)-1}\\ &\leq(2^{n}\beta\omega_{n}^{-1})^{g^{+}}\gamma F^{2}2^{g^{+}+1}\omega_{n}M^{f^{+}(R)-f^{-}(R)}R^{n-f^{+}(R)}|D_{t}|^{f^{-}(R)-1}\ \ (\text{since}\ 2^{-t}\psi\leq\psi\leq M)\\ &\leq(2^{n}\beta\omega_{n}^{-1})^{g^{+}}\gamma F^{2}2^{g^{+}+2}\omega_{n}R^{n-f^{+}(R)}|D_{t}|^{f^{-}(R)-1},\quad(t=0,\ldots,s-2).\end{split}$ (3.28)
Consequently, by inequality (3.28), we deduce that
$|A_{k_{s-1},\frac{R}{2}}|^{\frac{(n-1)f^{-}(R)}{n}}\leq c_{1}R^{n-f^{+}(R)}|D_{t}|^{f^{-}(R)-1},\quad(t=0,\ldots,s-2),$ (3.29)
where $c_{1}=(2^{n}\beta\omega_{n}^{-1})^{g^{+}}\gamma F^{2}2^{g^{+}+2}\omega_{n}+1=c(n,g^{+},F,\beta,\gamma)>1$.
Thus, due to inequality (3.29), we infer that
$|A_{k_{s-1},\frac{R}{2}}|^{\frac{(n-1)f^{-}(R)}{n(f^{-}(R)-1)}}\leq c_{2}R^{\frac{n-f^{+}(R)}{(f^{-}(R)-1)}}|D_{t}|,\quad(t=0,\ldots,s-2),$ (3.30)
where $c_{2}:=c_{1}^{\frac{1}{g^{-}-1}}=c(n,g^{-},g^{+},F,\beta,\gamma)$.
Summing inequality (3.30) with respect to $t$ from 0 to s2s-2 and noting that
$\sum_{t=0}^{s-2}|D_{t}|\leq|B_{\frac{R}{2}}|=\omega_{n}\left(\frac{R}{2}\right)^{n},$
we obtain
$|A_{k_{s-1},\frac{R}{2}}|^{\frac{(n-1)f^{-}(R)}{n(f^{-}(R)-1)}}\leq\frac{c_{2}2^{-n}\omega_{n}}{s-1}R^{\frac{n-f^{+}(R)}{(f^{-}(R)-1)}+n}=\frac{c_{2}2^{-n}\omega_{n}}{s-1}R^{\frac{nf^{-}(R)-f^{+}(R)}{(f^{-}(R)-1)}},$
and therefore, using Lemma 3.4, we get
$\begin{split}|A_{k_{s-1},\frac{R}{2}}|&\leq\left(\frac{c_{2}2^{-n}\omega_{n}}{s-1}\right)^{\frac{n(f^{-}(R)-1)}{(n-1)f^{-}(R)}}R^{\frac{nf^{-}(R)-f^{+}(R)}{(f^{-}(R)-1)}\times\frac{n(f^{-}(R)-1)}{(n-1)f^{-}(R)}}\\ &\leq\left(\frac{c_{2}2^{-n}\omega_{n}}{s-1}\right)^{\frac{n(f^{-}(R)-1)}{(n-1)f^{-}(R)}}R^{\frac{nf^{-}(R)-f^{+}(R)}{f^{-}(R)}\times\frac{n}{(n-1)}}\\ &\leq\left(\frac{c_{2}2^{-n}\omega_{n}}{s-1}\right)^{\frac{n(f^{-}(R)-1)}{(n-1)f^{-}(R)}}R^{\frac{n\left(f^{-}(R)-f^{+}(R)\right)}{f^{-}(R)}\times\frac{n}{(n-1)}}R^{n}\\ &\leq\left(\frac{c_{2}2^{-n}\omega_{n}}{s-1}\right)^{\frac{n(f^{-}(R)-1)}{(n-1)f^{-}(R)}}L^{\frac{n}{f^{-}(R)}\times\frac{n}{(n-1)}}R^{n}\\ &\leq\left(\frac{c_{2}2^{-n}\omega_{n}}{s-1}\right)^{\frac{n(f^{-}(R)-1)}{(n-1)f^{-}(R)}}L^{\frac{n}{g^{-}}\times\frac{n}{(n-1)}}R^{n}.\end{split}$ (3.31)
Now we choose a natural number $s$ such that $\frac{c_{2}2^{-n}\omega_{n}}{s-1}<1$ and
$\left(\frac{c_{2}2^{-n}\omega_{n}}{s-1}\right)^{\frac{n(g^{-}-1)}{(n-1)g^{-}}}L^{\frac{n}{g^{-}}\times\frac{n}{(n-1)}}\leq\theta.$ (3.32)
Note that $s=s(\theta,c_{2},n,g^{-},L)=s(\theta,n,g^{-},g^{+},F,L,\beta,\gamma)$. Then for such $s$, when inequality (3.20) does not hold, from inequalities (3.31) and (3.32), we have
$|A_{k_{s-1},\frac{R}{2}}|\leq\theta R^{n}$
which means $|A_{k^{0},\frac{R}{2}}|\leq\theta R^{n}$, i.e., inequality (3.21) holds, since
$k^{0}=\max_{B_{R}}w(x)-2^{-s+1}\psi=k_{s-1}.$
Lemma 3.9 is proved. ∎
Lemma 3.10.
There is a number $s=s(n,g^{-},g^{+},L,F,\gamma)>2$ such that
$\psi\leq 2^{s}\max\left\{\max_{B_{R}}w(x)-\max_{B_{\frac{R}{4}}}w(x),\left(\frac{\gamma+\gamma_{1}+1}{\gamma}\right)^{\frac{1}{f^{+}(R)}}R\right\},\text{ for any }R\in(0,R_{0}].$ (3.33)
Proof.
Let $R\in(0,R_{0}]$, and let
$\theta=\theta(n,g^{-},g^{+},F,L,\gamma)$
be the constant appearing in Lemma 3.8. Applying Lemma 3.9 with this choice of $\theta$, we obtain a constant
$s=s(\theta,n,g^{-},g^{+},F,L,\gamma)=s(n,g^{-},g^{+},F,L,\gamma)$
such that at least one of the inequalities (3.20) or (3.21) holds. If (3.20) holds, then (3.33) follows immediately. Now suppose that (3.20) does not hold. Then, by Lemma 3.9, inequality (3.21) must be satisfied; that is,
$|A_{k^{0},\frac{R}{2}}|\leq\theta R^{n},$
where $k^{0}=\max_{B_{R}}w(x)-2^{-s+1}\psi$. Set $H>2^{-s+1}\psi$. Then, since inequality (3.20) does not hold, we obtain
$H>2^{-s+1}\cdot 2^{s}\left(\frac{\gamma+\gamma_{1}+1}{\gamma}\right)^{\frac{1}{f^{+}(R)}}R>\left(\frac{\gamma+\gamma_{1}+1}{\gamma}\right)^{\frac{1}{f^{+}(R)}}R.$
Thus, by Lemma 3.8, one has
$\max_{B_{\frac{R}{4}}}w(x)\leq k^{0}+\frac{H}{2}=\max_{B_{R}}w(x)-2^{-s}\psi$
i.e.,
$\psi\leq 2^{s}\left(\max_{B_{R}}w(x)-\max_{B_{\frac{R}{4}}}w(x)\right).$
This shows that inequality (3.33) holds. Lemma 3.10 is proved. ∎
Lemma 3.11.
For any $R\in(0,R_{0}]$, at least one of the following two inequalities holds:
$\operatorname{osc}_{B_{R}}u\leq\tau 2^{s}\frac{\gamma+\gamma_{1}+1}{\gamma}R,$ (3.34)
or
$\operatorname{osc}_{B_{\frac{R}{4}}}u\leq(1-\tau^{-1}2^{-s})\operatorname{osc}_{B_{R}}u,$ (3.35)
where $\tau=\max\left\{2,\frac{2}{\delta}\right\}$, $s=s(n,g^{-},g^{+},F,L,\gamma)$ is the constant as in Lemma 3.10.
Proof.
By Lemma 3.10, inequality (3.33) holds, this implies that at least one of the following two inequalities holds:
$\psi\leq 2^{s}\left(\frac{\gamma+\gamma_{1}+1}{\gamma}\right)^{\frac{1}{f^{+}(R)}}R,$ (3.36)
or
$\psi\leq 2^{s}\left(\max_{B_{R}}w(x)-\max_{B_{\frac{R}{4}}}w(x)\right).$ (3.37)
When inequality (3.36) holds, we have
$\operatorname{osc}_{B_{R}}u=\tau\psi\leq\tau 2^{s}\left(\frac{\gamma+\gamma_{1}+1}{\gamma}\right)^{\frac{1}{f^{+}(R)}}R\leq\tau 2^{s}\frac{\gamma+\gamma_{1}+1}{\gamma}R,$
i.e., inequality (3.34) holds. When inequality (3.37) holds, we have
$\operatorname{osc}_{B_{R}}u,$ $=\tau\psi\leq\tau 2^{s}\left(\max_{B_{R}}w(x)-\max_{B_{\frac{R}{4}}}w(x)\right)$
$\leq\tau 2^{s}[\operatorname{osc}_{B_{R}}w-\operatorname{osc}_{B_{\frac{R}{4}}}w]$
$=\tau 2^{s}\left[\operatorname{osc}_{B_{R}}u-\operatorname{osc}_{B_{\frac{R}{4}}}u\right]$
this implies inequality (3.35). Lemma 3.11 is proved. ∎
We now return to the proof of Theorem 1.2. Combining the result of Lemma 3.3 with the conclusion of Lemma 3.11, we deduce that
$u\in C^{0,\alpha}(B_{R_{0}}(x_{0})),$
where
$\alpha=\min\left\{1,-\log_{4}(1-\tau^{-1}2^{-s})\right\}=\alpha(n,g^{-},g^{+},F,L,\gamma,\delta).$
By the arbitrarily of $x_{0}\in\Omega$, we have $u\in C^{0,\alpha}(\Omega)$. The proof of Theorem 1.2 is completed. ∎
Proof of Theorem 1.5.
Following the same lines as the proof of Theorem 1.2, we can establish the proof of Theorem 1.5. ∎
Remark 3.12.
In analogy with Section 7 of Chapter 2 in [49], one can define the class
$\mathcal{B}_{G(x,t)}(\Omega\cup S_{1},M,\gamma,\gamma_{1},\delta),$
where $S_{1}\subset\partial\Omega$, and establish the corresponding results.
4. Application to Problems in Musielak–Orlicz–Sobolev Spaces
This section is devoted to the regularity of weak solutions to problem ($\mathcal{P}$), with the ultimate goal of proving Theorems 1.13 and 1.14.
A fundamental step in establishing Hölder continuity ($C^{0,\alpha}$) is to first prove that weak solutions are bounded. Leveraging the embedding results of Diening and Cianchi [14], we begin by demonstrating the boundedness of weak solutions to ($\mathcal{P}$) under both Dirichlet and Neumann boundary conditions, encompassing both critical and subcritical growth conditions on the nonlinearity. This is the subject of Theorems 1.9 and 1.10.
Our strategy for proving boundedness is inspired by the work of Ho, Kim, Winkert, and Zhang [43]. We extend their approach to provide a unified treatment for both critical and subcritical cases, thereby yielding a more general result.
4.1. Boundedness of Weak Solutions to Problems with Dirichlet Boundary Condition
In this subsection, we focus on proving the boundedness of weak solutions to the Dirichlet problem ($\mathcal{PD}$), as stated in Theorem 1.9. To this end, we need the following lemma concerning the geometric convergence of sequences of numbers will be needed for the De Giorgi iteration arguments below. It can be found in Ho-Sim [44, Lemma 4.3]. The case $\mu_{1}=\mu_{2}$ is contained in Ladyženskaja-Solonnikov-Uralceva [49, Chapter II, Lemma 5.6], see also DiBenedetto [22, Chapter I, Lemma 4.1].
Lemma 4.1.
Let $\{Z_{h}\}_{h\geq 0}$ be a sequence of positive numbers satisfying the recursion inequality
$Z_{h+1}\leq Kb^{h}\left(Z_{h}^{1+\mu_{1}}+Z_{h}^{1+\mu_{2}}\right),\quad h=0,1,2,\dots,$
for some constants $K>0$, $b>1$, and exponents $\mu_{2}\geq\mu_{1}>0$. Suppose that either
$Z_{0}\leq\min\left(1,(2K)^{-\frac{1}{\mu_{1}}}b^{-\frac{1}{\mu_{1}^{2}}}\right),$
or
$Z_{0}\leq\min\left((2K)^{-\frac{1}{\mu_{1}}}b^{-\frac{1}{\mu_{1}^{2}}},(2K)^{-\frac{1}{\mu_{2}}}b^{-\frac{1}{\mu_{1}\mu_{2}}-\frac{\mu_{2}-\mu_{1}}{\mu_{2}^{2}}}\right).$
Then there exists an index $h_{0}=0,1,2,\ldots$ such that $Z_{h}\leq 1$ for all $h\geq h_{0}$, and moreover:
$Z_{h}\leq\min\left(1,(2K)^{-\frac{1}{\mu_{1}}}b^{-\frac{1}{\mu_{1}^{2}}}b^{-\frac{h}{\mu_{1}}}\right),\quad\forall h\geq h_{0}.$
In particular, $Z_{h}\to 0$ as $h\to\infty$.
Proof of Theorem 1.9.
The compactness of $\overline{\Omega}$ implies that for any $R>0$, there exists a finite open cover $\{B_{i}(R)\}_{i=1}^{m}$ of balls with radius $R$, such that
$\overline{\Omega}\subset\bigcup_{i=1}^{m}B_{i},$
where each $\Omega_{i}:=B_{i}\cap\Omega$ (for $i=1,\ldots,m$) is also a Lipschitz domain (see, for example, Papageorgiou–Winkert [77, Theorem 1.4.86]).
Let $u$ be a weak solution of problem ($\mathcal{PD}$). Let $k_{*}\geq 1$ be sufficiently large such that
$\int_{A_{k_{*}}}G(x,|\nabla u|){\rm d}x+\int_{A_{k_{*}}}H_{1}(x,|u|){\rm d}x<1,$ (4.1)
where $A_{k}:=\{x\in\Omega\,:\,u(x)>k\}$ for $k\in\mathbb{R}$.
We define
$Z_{h}:=\int_{A_{k_{h}}}G(x,|\nabla u|){\rm d}x+\int_{A_{k_{h}}}H_{1}(x,(u-k_{h})){\rm d}x,$ (4.2)
where
$k_{h}:=k_{*}\left(2-\frac{1}{2^{h}}\right)\quad\text{for }h=0,1,2,\ldots.$
Obviously, $\text{for }h=0,1,2,\ldots$, it holds
$k_{h}\nearrow 2k_{*}\quad\text{and}\quad k_{*}\leq k_{h}<2k_{*}.$ (4.3)
Since $k_{h}<k_{h+1}$ and $A_{k_{h+1}}\subset A_{k_{h}}$ for all $h=0,1,2,\ldots$, we have
$Z_{h+1}\leq Z_{h}\quad\text{for all }h=0,1,2,\ldots.$ (4.4)
Moreover, for $x\in A_{k_{h+1}}$, we see that
$u(x)-k_{h}\geq u(x)\left(1-\frac{k_{h}}{k_{h+1}}\right)=\frac{u(x)}{2^{h+2}-1}.$
Hence, we obtain
$u(x)\leq(2^{h+2}-1)(u(x)-k_{h})\quad\text{for a.\,a.\,}x\in A_{k_{h+1}}\text{ and for all }h=0,1,2,\ldots.$ (4.5)
Furthermore, by Propositions 2.11 and 2.16, we have
$\begin{split}|A_{k_{h+1}}|&\leq\int_{A_{k_{h+1}}}H_{1}\left(x,\frac{u-k_{h}}{k_{h+1}-k_{h}}\right)dx\\ &\leq\max\left\{\frac{1}{(k_{h+1}-k_{h})^{h^{-}_{1}}},\frac{1}{(k_{h+1}-k_{h})^{h^{+}_{1}}}\right\}\int_{A_{k_{h}}}H_{1}\left(x,u-k_{h}\right)dx\\ &\leq\max\left\{\frac{2^{h^{-}_{1}(h+1)}}{k_{*}^{h^{-}}},\frac{2^{h^{+}(h+1)}}{k_{*}^{h^{+}_{1}}}\right\}\int_{A_{k_{h}}}H_{1}\left(x,u-k_{h}\right)dx\\ &\leq\frac{2^{h^{+}_{1}(h+1)}}{k_{*}^{h^{-}_{1}}}\int_{A_{k_{h}}}H_{1}\left(x,u-k_{h}\right)dx\\ &\leq\frac{2^{h^{+}_{1}(h+1)}}{k_{*}^{h^{-}_{1}}}Z_{h}\\ &\leq 2^{h^{+}_{1}(h+1)}Z_{h}\quad\text{for all }h=0,1,2,\ldots.\end{split}$ (4.6)
In the remainder of the proof, we denote by $C_{i}$ (for $i=1,2,\ldots$) positive constants that are independent of $h$ and $k_{*}$.
Claim 1: There exist positive constants $\mu_{1}$ and $\mu_{2}$ such that
$\int_{A_{k_{h+1}}}H_{1}(x,(u-k_{h+1})){\rm d}x\leq C_{1}2^{\frac{h\left(h^{+}_{1}\right)^{2}}{g^{-}}}\left(Z_{h}^{1+\mu_{1}}+Z_{h}^{1+\mu_{2}}\right)\quad\text{for all }h=0,1,2,\ldots.$
First, note that
$\begin{split}\int_{A_{k_{h+1}}}H_{1}(x,(u-k_{h+1})){\rm d}x&=\int_{\Omega}H_{1}(x,(u-k_{h+1})^{+}){\rm d}x\\ &\leq\sum_{i=1}^{m}\int_{\Omega_{i}}H_{1}(x,(u-k_{h+1})^{+}){\rm d}x.\end{split}$ (4.7)
Let $i\in\{1,\ldots,m\}$. From (4.1), Propositions 2.11, 2.16 and 2.17 for $\Omega=\Omega_{i}$ we have
$\int_{\Omega_{i}}H_{1}(x,(u-k_{h+1})^{+}){\rm d}x$ $\leq\|(u-k_{h+1})^{+}\|_{L^{H_{1}(x,t)}(\Omega_{i})}^{h^{-}_{1}}$
$\leq C_{2}\left[\|\nabla(u-k_{h+1})^{+}\|_{L^{G(x,t)}(\Omega_{i})}+\|(u-k_{h+1})^{+}\|_{L^{G(x,t)}(\Omega_{i})}\right]^{h^{-}_{1}}.$
Then, again by Propositions 2.11, 2.16 and 2.17 along with (4.1), this leads to
$\int_{\Omega_{i}}H_{1}(x,(u-k_{h+1})^{+}){\rm d}x$ $\leq C_{3}\left(\int_{\Omega_{i}}G(x,|\nabla(u-k_{h+1})^{+}|){\rm d}x+\int_{\Omega_{i}}G(x,(u-k_{h+1})^{+})dx\right)^{\frac{h^{-}}{g^{+}}}$
$\leq C_{3}\left(\int_{A_{k_{h+1}}}G(x,|\nabla u|)dx+\int_{A_{k_{h+1}}}H_{1}(x,(u-k_{h+1})){\rm d}x+|A_{k_{h+1}}|\right)^{\frac{h^{-}_{1}}{g^{+}}}.$
From this, (4.2), (4.4) and (4.6), we obtain
$\int_{\Omega_{i}}H_{1}(x,(u-k_{h+1})^{+}){\rm d}x\leq C_{4}2^{\frac{hh^{+}_{1}h^{-}_{1}}{g^{+}}}Z_{h}^{\frac{h^{-}_{1}}{g^{+}}}.$
Combining this with (4.7), we deduce that
$\int_{\Omega}H_{1}(x,(u-k_{h+1})^{+}{\rm d}x\leq C_{5}2^{\frac{h\left(h^{+}_{1}\right)^{2}}{g^{-}}}\left(Z_{h}^{1+\mu_{1}}+Z_{h}^{1+\mu_{2}}\right),$
where
$0<\mu_{1}:=\frac{h^{-}_{1}}{g^{+}}-1=:\mu_{2}.$
This proves Claim 1.
Claim 2: It holds that
$\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x\leq C_{6}2^{h\left[\frac{\left(h^{+}\right)^{2}}{g^{-}}+h^{+}\right]}\left(Z_{h-1}^{1+\mu_{1}}+Z_{h-1}^{1+\mu_{2}}\right)\quad\text{for all }h=1,2,\ldots.$
Testing ($\mathcal{FDV}$) with $\varphi=(u-k_{h+1})^{+}\in W_{0}^{1,G(x,t)}(\Omega)$, it yields
$\int_{\Omega}A(x,u,\nabla u)\cdot\nabla(u-k_{h+1})^{+}dx=\int_{\Omega}B(x,u,\nabla u)(u-k_{h+1})^{+}{\rm d}x,$
which can be written as
$\int_{A_{k_{h+1}}}A(x,u,\nabla u)\cdot\nabla udx=\int_{A_{k_{h+1}}}B(x,u,\nabla u)(u-k_{h+1})dx.$ (4.8)
We proceed to estimate the terms in (4.8), starting with the left-hand side.
Since $u\geq u-k_{h+1}>0$, and $u>k_{h+1}\geq 1$ on the set $A_{k_{h+1}}$. Furthermore, the function $t\mapsto H_{1}(x,t)$ is increasing for all $x\in\Omega$. Therefore, by assumption ($\mathcal{A}_{1}$) and Propositions 2.11 and 2.16, we obtain
$\int_{A_{k_{h+1}}}A(x,u,\nabla u)\cdot\nabla u\,\mathrm{d}x$ $\geq a_{4}\int_{A_{k_{h+1}}}G(x,|\nabla u|)\,\mathrm{d}x-a_{5}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x-\int_{A_{k_{h+1}}}a_{6}\,\mathrm{d}x$
$\geq a_{4}\int_{A_{k_{h+1}}}G(x,|\nabla u|)\,\mathrm{d}x-\max\{a_{5},a_{6}\}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x.$
We now estimate the right-hand side of (4.8). Let $\varepsilon\in(0,1]$ be chosen such that
$b_{1}\varepsilon^{\frac{h^{+}_{1}}{h^{+}_{1}-1}}\leq\frac{a_{4}}{4}.$
Then, employing assumption ($\mathcal{B}$), Young’s inequality (2.3), the estimate (2.4), and Propositions 2.11 and 2.16, we find
$\int_{A_{k_{h+1}}}B(x,u,\nabla u)(u-k_{h+1})\,\mathrm{d}x$ $\leq b_{1}\int_{A_{k_{h+1}}}M(x,|\nabla u|)u\,\mathrm{d}x+b_{2}\int_{A_{k_{h+1}}}h_{1}(x,u)u\,\mathrm{d}x+\int_{A_{k_{h+1}}}b_{3}(u-k_{h+1})\,\mathrm{d}x$
$\leq b_{1}\int_{A_{k_{h+1}}}\frac{\varepsilon}{\varepsilon}M(x,|\nabla u|)u\,\mathrm{d}x+\max\{h^{+}_{1}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x$
$\leq b_{1}\int_{A_{k_{h+1}}}\widetilde{H_{1}}\left(x,\varepsilon M(x,|\nabla u|)\right)\,\mathrm{d}x+b_{1}\int_{A_{k_{h+1}}}H_{1}\left(x,\frac{1}{\varepsilon}u\right)\,\mathrm{d}x$
$\quad+\max\{h^{+}_{1}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x$
$\leq b_{1}\varepsilon^{\frac{h^{+}_{1}}{h^{+}_{1}-1}}\int_{A_{k_{h+1}}}\widetilde{H_{1}}\left(x,M(x,|\nabla u|)\right)\,\mathrm{d}x+b_{1}\varepsilon^{-h^{+}_{1}}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x$
$\quad+\max\{h^{+}_{1}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x$
$\leq b_{1}\varepsilon^{\frac{h^{+}_{1}}{h^{+}_{1}-1}}\int_{A_{k_{h+1}}}G(x,|\nabla u|)\,\mathrm{d}x+C_{7}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x,$
where
$C_{7}:=\max\left\{h^{+}_{1}b_{2},b_{3},b_{1}\varepsilon^{-h^{+}_{1}}\right\}$.
Using the choice of $\varepsilon$, we conclude that
$\int_{A_{k_{h+1}}}B(x,u,\nabla u)(u-k_{h+1})\,\mathrm{d}x\leq\frac{a_{4}}{4}\int_{A_{k_{h+1}}}G(x,|\nabla u|)\,\mathrm{d}x+C_{7}\int_{A_{k_{h+1}}}H_{1}(x,u)\,\mathrm{d}x.$
Combining the estimates above with (4.8) and then using (4.5), we obtain
$\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x$ $\leq C_{8}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x$
$\leq C_{8}\int_{A_{k_{h+1}}}H_{1}\left(x,(2^{h+2}-1)(u-k_{h})\right){\rm d}x.$
Hence,
$\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x\leq C_{9}2^{hh^{+}_{1}}\int_{\Omega}H_{1}(x,(u-k_{h})^{+}){\rm d}x.$
Then, Claim 2 follows from the preceding inequality together with Claim 1.
From Claims 1 and 2, along with (4.4), we conclude that
$Z_{h+1}\leq C_{10}b^{h}\left(Z_{h-1}^{1+\mu_{1}}+Z_{h-1}^{1+\mu_{2}}\right)\quad\text{for all }h=1,2,\ldots,$ (4.9)
where
$b:=2^{\big[\frac{\left(h^{+}_{1}\right)^{2}}{g^{-}}+h^{+}_{1}\big]}>1.$
This yields to
$Z_{2(h+1)}\leq C_{10}b^{2h+1}\left(Z_{2h}^{1+\mu_{1}}+Z_{2h}^{1+\mu_{2}}\right)\quad\text{for all }h=0,1,2,\ldots,$
that is,
$\tilde{Z}_{h+1}\leq bC_{10}\tilde{b}^{h}\left(\tilde{Z}_{h}^{1+\mu_{1}}+\tilde{Z}_{h}^{1+\mu_{2}}\right)\quad\text{for all }h=0,1,2,\ldots,$ (4.10)
where $\tilde{Z}_{h}:=Z_{2h}$ and $\tilde{b}:=b^{2}$. Applying Lemma 4.1 to (4.10) yields
$Z_{2h}=\tilde{Z}_{h}\to 0\quad\text{as }h\to\infty$ (4.11)
provided that
$\tilde{Z}_{0}\leq\min\left\{(2bC_{10})^{-\frac{1}{\mu_{1}}}\ \tilde{b}^{-\frac{1}{\mu_{1}^{2}}},\left(2bC_{10}\right)^{-\frac{1}{\mu_{2}}}\ \tilde{b}^{-\frac{1}{\mu_{1}\mu_{2}}-\frac{\mu_{2}-\mu_{1}}{\mu_{2}^{2}}}\right\}.$ (4.12)
Again, from (4.9) we obtain
$Z_{2(h+1)+1}\leq C_{10}b^{2(h+1)}\left(Z_{2h+1}^{1+\mu_{1}}+Z_{2h+1}^{1+\mu_{2}}\right)\quad\text{for all }h=0,1,2,\ldots,$
which can be written as
$\hat{Z}_{h+1}\leq\tilde{b}C_{10}\tilde{b}^{h}\left(\hat{Z}_{h}^{1+\mu_{1}}+\hat{Z}_{h}^{1+\mu_{2}}\right)\quad\text{for all }h=1,2,\ldots,$ (4.13)
where $\hat{Z}_{h}:=Z_{2h+1}$. From Lemma 4.1 applied to (4.13) it follows that
$Z_{2h+1}=\hat{Z}_{h}\to 0\quad\text{as }n\to\infty$ (4.14)
provided that
$\hat{Z}_{0}\leq\min\left\{(2\tilde{b}C_{10})^{-\frac{1}{\mu_{1}}}\ \tilde{b}^{-\frac{1}{\mu_{1}^{2}}},\left(2\tilde{b}C_{10}\right)^{-\frac{1}{\mu_{2}}}\ \tilde{b}^{-\frac{1}{\mu_{1}\mu_{2}}-\frac{\mu_{2}-\mu_{1}}{\mu_{2}^{2}}}\right\}.$ (4.15)
Note that
$\hat{Z}_{0}=Z_{1}\leq Z_{0}=\tilde{Z}_{0}\leq\int_{A_{k_{*}}}G(x,|\nabla u|)dx+\int_{A_{k_{*}}}H(x,u){\rm d}x.$
Therefore, by choosing $k_{*}>1$ sufficiently large we have
$\int_{A_{k_{*}}}G(x,|\nabla u|)dx+\int_{A_{k_{*}}}H_{1}(x,u){\rm d}x$ $\leq\min\left\{1,(2\tilde{b}C_{10})^{-\frac{1}{\mu_{1}}}\ \tilde{b}^{-\frac{1}{\mu_{1}^{2}}},\left(2\tilde{b}C_{10}\right)^{-\frac{1}{\mu_{2}}}\ \tilde{b}^{-\frac{1}{\mu_{1}\mu_{2}}-\frac{\mu_{2}-\mu_{1}}{\mu_{2}^{2}}}\right\}.$
Hence, (4.1), (4.12) and (4.15) are fulfilled and we obtain (4.11) and (4.14). This means that
$Z_{h}=\int_{A_{k_{h}}}G(x,|\nabla u|)dx+\int_{A_{k_{h}}}H_{1}(x,(u-k_{h}))dx\to 0\quad\text{as }h\to\infty.$
In particular, we have
$\int_{\Omega}H_{1}(x,(u-2k_{*})^{+})dx=0.$
Consequently, by Proposition 2.11, $(u-2k_{*})^{+}=0$ a. e. in $\Omega$ and so
$\underset{\Omega}{\mathop{\rm ess\,sup}}\ u\leq 2k_{*}.$
Replacing $u$ by u-u in the arguments above, we also obtain
$\underset{\Omega}{\mathop{\rm ess\,sup}}\ (-u)\leq 2k_{*}.$
Hence, $\|u\|_{\infty}\leq 2k_{*}$. This finishes the proof. ∎
4.2. Boundedness of Weak Solutions to Problems with Neumann Boundary Condition
In this subsection, we address the boundedness of weak solutions to the Neumann problem ($\mathcal{PN}$), as formulated in Theorem 1.10.
Proof of Theorem 1.10.
As before, since $\overline{\Omega}$ is compact, for any $R>0$, there exists a finite open cover $\{B_{i}(R)\}_{i=1}^{m}$ of balls $B_{i}:=B_{i}(R)$, each of radius $R$, such that
$\overline{\Omega}\subset\bigcup_{i=1}^{m}B_{i},$
and each intersection $\Omega_{i}:=B_{i}\cap\Omega$ (for $i=1,\ldots,m$) is also a Lipschitz domain.
We denote by $I\subset\{1,\ldots,m\}$ the set of indices for which $B_{i}\cap\partial\Omega\neq\emptyset$, and we choose $R>0$ sufficiently small.
Let $u$ be a weak solution to problem ($\mathcal{PN}$), and let $k_{*}\geq 1$ be sufficiently large such that
$\int_{A_{k_{*}}}G(x,|\nabla u|){\rm d}x+\int_{A_{k_{*}}}H_{1}(x,u){\rm d}x+\int_{\Gamma_{k_{*}}}H_{2}(x,u){\rm d}\sigma<1,$ (4.16)
where
$A_{k}:=\{x\in\Omega\,:\,u(x)>k\}\quad\text{and}\quad\Gamma_{k}:=\{x\in\partial\Omega\,:\,u(x)>k\}\quad\text{for }k\in\mathbb{R}.$
We define
$Z_{h}:=\int_{A_{k_{h}}}G(x,|\nabla u|){\rm d}x+\int_{A_{k_{h}}}H_{1}\left(x,(u-k_{h})\right){\rm d}x+\int_{\Gamma_{k_{n}}}H_{2}\left(x,(u-k_{h})\right){\rm d}\sigma,$ (4.17)
where $\{k_{h}\}_{h\geq 0}$ is a sequence defined as in the proof of Theorem 1.9. Since $k_{h}<k_{h+1}$, we have $A_{k_{h+1}}\subset A_{k_{h}}$ and $\Gamma_{k_{h+1}}\subset\Gamma_{k_{h}}$ for all $h=0,1,2,\ldots$. Hence, it holds
$Z_{h+1}\leq Z_{h}\quad\text{for all }h=0,1,2,\ldots.$ (4.18)
Moreover, we have the following estimates (see the proof of Theorem 1.9)
$u(x)$ $\leq\left(2^{h+2}-1\right)(u(x)-k_{h})\quad\text{for a.\,a.\,}x\in A_{k_{h+1}}\text{ and for all }h=0,1,2,\ldots,$ (4.19)
$u(x)$ $\leq\left(2^{h+2}-1\right)(u(x)-k_{h})\quad\text{for a.\,a.\,}x\in\Gamma_{k_{h+1}}\text{ and for all }h=0,1,2,\ldots.$ (4.20)
As shown in (4.6) the following estimate holds true
$|A_{k_{h+1}}|\leq\frac{2^{h_{1}^{+}(h+1)}}{k_{*}^{h^{-}_{1}}}Z_{h}\leq 2^{h^{+}_{1}(h+1)}Z_{h}\quad\text{for all }h=0,1,2,\ldots.$ (4.21)
In the following, we will denote by $C_{i}$ ($i=1,2,3,\ldots$) positive constants independent of $h$ and $k_{*}$.
Claim 1: It holds that
$\int_{A_{k_{h+1}}}H_{1}\left(x,(u-k_{h+1})\right){\rm d}x\leq C_{1}2^{\frac{h\left(h^{+}_{1}\right)^{2}}{g^{-}}}\left(Z_{h}^{1+\nu_{1}}+Z_{h}^{1+\nu_{2}}\right)\quad\text{for all }h=0,1,2,\ldots,$
where
$0<\nu_{1}:=\frac{h_{1}^{-}}{g^{+}}-1=:\nu_{2}.$
The proof of Claim 1 is the same as Claim 1 in Theorem 1.9.
Claim 2: There exist positive constants $\nu_{3}$ and $\nu_{4}$ such that
$\int_{\Gamma_{k_{h+1}}}H_{2}\left(x,(u-k_{h+1})\right){\rm d}\sigma\leq C_{2}2^{\frac{h\left(h^{+}_{2}\right)^{2}}{g^{-}}}\left(Z_{h}^{1+\nu_{3}}+Z_{h}^{1+\nu_{4}}\right)\quad\text{for all }h=0,1,2,\ldots.$
First, we see that
$\begin{split}\int_{\Gamma_{k_{h+1}}}H_{2}(x,(u-k_{h+1})){\rm d}\sigma&=\int_{\partial\Omega}H_{2}(x,(u-k_{h+1})^{+}){\rm d}\sigma\\ &\leq\sum_{i=1}^{m}\int_{\partial\Omega\cap\partial\Omega_{i}}H_{2}(x,(u-k_{h+1})^{+}){\rm d}\sigma.\end{split}$ (4.22)
Let $i\in\{1,\ldots,m\}$. From Propositions 2.11 and 2.16, inequality (4.16), and Proposition 2.17, for $\Omega=\Omega_{i}$, we have
$\int_{\partial\Omega\cap\partial\Omega_{i}}H_{2}(x,(u-k_{h+1})^{+}){\rm d}\sigma$
$\leq\|(u-k_{h+1})^{+}\|_{L^{H_{2}(x,t)}(\partial\Omega\cap\partial\Omega_{i})}^{h_{2}^{-}}$
$\leq\|(u-k_{h+1})^{+}\|_{L^{H_{2}(x,t)}(\partial\Omega_{i})}^{h_{2}^{-}}$
$\leq C_{3}\left[\|\nabla(u-k_{n+1})^{+}\|_{L^{G(x,t)}(\Omega_{i})}+\|(u-k_{h+1})^{+}\|_{L^{G(x,t)}(\Omega_{i})}\right]^{h_{2}^{-}}.$
Then, by Propositions 2.11, 2.16 and inequality(4.16) it follows
$\int_{\partial\Omega\cap\partial\Omega_{i}}H_{2}(x,(u-k_{h+1})^{+}){\rm d}\sigma$
$\leq C_{4}\left(\int_{\Omega_{i}}G(x,|\nabla(u-k_{h+1})^{+}|){\rm d}x+\int_{\Omega_{i}}G(x,(u-k_{h+1})^{+}){\rm d}x\right)^{\frac{h_{2}^{-}}{g^{+}}}$
$\leq C_{4}\left(\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x+\int_{A_{k_{h+1}}}H_{2}(x,(u-k_{h+1})){\rm d}x+|A_{k_{h+1}}|\right)^{\frac{h_{2}^{-}}{g^{+}}}.$
From this combined with (4.17), (4.18) as well as (4.21) we obtain
$\int_{\partial\Omega\cap\partial\Omega_{i}}H_{2}(x,(u-k_{h+1})^{+}){\rm d}\sigma$ $\leq C_{5}2^{\frac{hh_{2}^{+}h_{2}^{-}}{g^{+}}}Z_{h}^{\frac{h_{2}^{-}}{g^{+}}}.$ (4.23)
Consequently, by (4.22) and (4.23), we conclude that
$\int_{\partial\Omega}H_{2}((u-k_{h+1})^{+}){\rm d}\sigma\leq C_{6}2^{\frac{h\left(h_{2}^{+}\right)^{2}}{g^{-}}}\left(Z_{h}^{1+\nu_{3}}+Z_{h}^{1+\nu_{4}}\right),$
where
$0<\nu_{3}:=\frac{h_{2}^{-}}{g^{+}}-1=:\nu_{4}.$
This proves Claim 2.
Claim 3: It holds that
$\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x\leq C_{7}2^{h\left[\frac{\left(h_{1}^{+}+h_{2}^{+}\right)^{2}}{g^{-}}+g_{*}^{+}\right]}\left(Z_{h-1}^{1+\mu_{1}}+Z_{h-1}^{1+\mu_{2}}\right)\quad\text{for all }h=1,2,3,\ldots,$
where
$0<\mu_{1}:=\nu_{1}=\nu_{2}=\nu_{3}=\nu_{4}=:\mu_{2}.$
Taking $\varphi=(u-k_{h+1})^{+}\in W^{1,G(x,t)}(\Omega)$ as test function in ($\mathcal{FNV}$) gives
$\int_{\Omega}A(x,u,\nabla u)\cdot\nabla\varphi{\rm d}x=\int_{\Omega}B(x,u,\nabla u)\varphi{\rm d}x+\int_{\partial\Omega}C(x,u)\varphi{\rm d}\sigma,$
which can be written as
$\begin{split}&\int_{A_{k_{h+1}}}A(x,u,\nabla u)\cdot\nabla u{\rm d}x\\ &=\int_{A_{k_{h+1}}}B(x,u,\nabla u)(u-k_{h+1}){\rm d}x+\int_{\Gamma_{k_{h+1}}}C(x,u)(u-k_{h+1}){\rm d}\sigma.\end{split}$ (4.24)
As done in the proof of Theorem 1.9, we apply assumptions ($\mathcal{A}_{1}$), ($\mathcal{B}$) and ($\mathcal{C}$), as well as Young’s inequality (2.3), (2.4) and Propositions 2.11 and 2.16, and we take $\varepsilon\in(0,1]$ such that
$b_{1}\varepsilon^{\frac{h_{1}^{+}}{h_{1}^{+}-1}}\leq\frac{a_{4}}{4},$
we have the estimates
$\begin{split}&\int_{A_{k_{h+1}}}A(x,u,\nabla u)\cdot\nabla u{\rm d}x\\ &\geq a_{4}\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x-\max\{a_{5},a_{6}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x,\end{split}$ (4.25)
and
$\begin{split}&\int_{A_{k_{h+1}}}B(x,u,\nabla u)(u-k_{h+1}){\rm d}x\\ &\leq b_{1}\int_{A_{k_{h+1}}}M(x,|\nabla u|)(u-k_{h+1}){\rm d}x+b_{2}\int_{A_{k_{h+1}}}h_{1}(x,u)(u-k_{h+1}){\rm d}x+b_{3}\int_{A_{k_{h+1}}}(u-k_{h+1})dx\\ &\leq b_{1}\int_{A_{k_{h+1}}}M(x,|\nabla u|)u{\rm d}x+b_{2}\int_{A_{k_{h+1}}}h_{1}(x,u)u{\rm d}x+b_{3}\int_{A_{k_{h+1}}}u(x)dx\\ &\leq b_{1}\int_{A_{k_{h+1}}}M(x,|\nabla u|)u{\rm d}x+b_{2}h_{1}^{+}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x+b_{3}\int_{A_{k_{h+1}}}H_{1}(x,u)dx\\ &\leq b_{1}\int_{A_{k_{h+1}}}\frac{\varepsilon}{\varepsilon}M(x,|\nabla u|)u{\rm d}x+\max\{h_{1}^{+}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x\\ &\leq b_{1}\int_{A_{k_{h+1}}}\widetilde{H_{1}}\left(x,\varepsilon M(x,|\nabla u|)\right){\rm d}x+b_{1}\int_{A_{k_{h+1}}}H_{1}\left(x,\frac{1}{\varepsilon}u\right){\rm d}x\\ &+\max\{h_{1}^{+}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x\\ &\leq b_{1}\varepsilon^{\frac{h_{1}^{+}}{h_{1}^{+}-1}}\int_{A_{k_{h+1}}}\widetilde{H_{1}}\left(x,M(x,|\nabla u|)\right){\rm d}x+b_{1}\varepsilon^{-h_{1}^{+}}\int_{A_{k_{h+1}}}H_{1}\left(x,u\right){\rm d}x\\ &+\max\{h_{1}^{+}b_{2},b_{3}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x\\ &\leq b_{1}\varepsilon^{\frac{h_{1}^{+}}{h_{1}^{+}-1}}\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x+\max\{h_{1}^{+}b_{2},b_{3},b_{1}\varepsilon^{-h_{1}^{+}}\}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x\\ &=\frac{a_{4}}{4}\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x+C_{7}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x.\end{split}$ (4.26)
Note that $u\geq u-k_{h+1}>0$ and $u>k_{h+1}\geq 1$ on $\Gamma_{k_{h+1}}$. So, applying assumption ($\mathcal{C}$), it yields that
$\begin{split}\int_{\Gamma_{k_{h+1}}}C(x,u)(u-k_{h+1}){\rm d}\sigma&\leq c_{1}\int_{\Gamma_{k_{h+1}}}h_{2}(x,u)(u-k_{h+1}){\rm d}\sigma+c_{2}\int_{\Gamma_{k_{h+1}}}(u-k_{h+1}){\rm d}\sigma\\ &\leq c_{1}\int_{\Gamma_{k_{h+1}}}h_{2}(x,u)u{\rm d}\sigma+c_{2}\int_{\Gamma_{k_{h+1}}}u{\rm d}\sigma\\ &\leq c_{1}h_{2}^{+}\int_{\Gamma_{k_{h+1}}}H_{2}(x,u){\rm d}\sigma+c_{2}\int_{\Gamma_{k_{h+1}}}H_{2}(x,u){\rm d}\sigma\\ &\leq(c_{1}h_{2}^{+}+c_{2})\int_{\Gamma_{k_{h+1}}}H_{2}(x,u){\rm d}\sigma.\end{split}$ (4.27)
Combining (4.24), (4.25), (4.26) and (4.27), and using (4.19) as well as (4.20), we obtain
$\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x$
$\leq C_{8}\int_{A_{k_{h+1}}}H_{1}(x,u){\rm d}x+C_{8}\int_{\Gamma_{k_{h+1}}}H_{2}(x,u){\rm d}\sigma$
$\leq C_{8}\int_{A_{k_{h+1}}}H_{1}\left(x,\left(2^{h+2}-1\right)(u-k_{h})\right){\rm d}x$
$\qquad+C_{8}\int_{\Gamma_{k_{h+1}}}H_{2}\left(x,\left(2^{h+2}-1\right)(u-k_{h})\right){\rm d}\sigma.$
Hence,
$\int_{A_{k_{h+1}}}G(x,|\nabla u|){\rm d}x\leq C_{9}2^{h(h_{1}^{+}+h_{2}^{+})}\left[\int_{A_{k_{h}}}H_{1}(x,(u-k_{h})){\rm d}x+\int_{\Gamma_{k_{h}}}H_{2}(x,(u-k_{h})){\rm d}\sigma\right].$
Then, Claim 3 follows from the last inequality and Claims 1 and 2.
From Claims 1, 2, and 3, along with (4.18) one has
$Z_{h+1}\leq C_{10}b^{h}\left(Z_{h-1}^{1+\mu_{1}}+Z_{h-1}^{1+\mu_{2}}\right)\quad\text{for all }h=1,2,3,\ldots,$ (4.28)
where
$b:=2^{\left[\frac{\left(h_{1}^{+}+h_{2}^{+}\right)^{2}}{g^{-}}+g_{*}^{+}\right]}>1.$
Repeating the same arguments used in the proof of Theorem 1.9, by choosing $k_{*}>1$ sufficiently large such that
$\int_{A_{k_{*}}}G(x,|\nabla u|){\rm d}x+\int_{A_{k_{*}}}H_{1}(x,u){\rm d}x+\int_{\Gamma_{k_{*}}}H_{2}(x,u){\rm d}\sigma$
$\leq\min\left\{1,(2\tilde{b}C_{10})^{-\frac{1}{\mu_{1}}}\ \tilde{b}^{-\frac{1}{\mu_{1}^{2}}},\left(2\tilde{b}C_{10}\right)^{-\frac{1}{\mu_{2}}}\ \tilde{b}^{-\frac{1}{\mu_{1}\mu_{2}}-\frac{\mu_{2}-\mu_{1}}{\mu_{2}^{2}}}\right\},$
where $\tilde{b}:=b^{2}$, we deduce from (4.28) that
$Z_{h}=\int_{A_{k_{h}}}G(x,|\nabla u|){\rm d}x+\int_{A_{k_{h}}}H_{1}(x,(u-k_{h})){\rm d}x+\int_{\Gamma_{k_{h}}}H_{2}(x,(u-k_{h})){\rm d}\sigma\to 0$
as $h\to\infty$. This implies that
$\int_{\Omega}H_{1}(x,(u-2k_{*})^{+}){\rm d}x+\int_{\partial\Omega}H_{2}(x,(u-2k_{*})^{+}){\rm d}\sigma=0.$
Therefore, $(u-2k_{*})^{+}=0$ a. e. in $\Omega$ and $(u-2k_{*})^{+}=0$ a. e. on $\partial\Omega$. This means that
$\underset{\Omega}{\mathop{\rm ess\,sup}}\ u+\underset{\partial\Omega}{\mathop{\rm ess\,sup}}\ u\leq 4k_{*}.$
By replacing $u$ with u-u in the arguments above, we can show in the same way that
$\underset{\Omega}{\mathop{\rm ess\,sup}}(-u)+\underset{\partial\Omega}{\mathop{\rm ess\,sup}}(-u)\leq 4k_{*}.$
Hence,
$\|u\|_{L^{\infty}(\Omega)}+\|u\|_{L^{\infty}(\partial\Omega)}\leq 4k_{*}.$
The proof is complete. ∎
4.3. Hölder Continuity of Weak Solutions to Problems with Dirichlet or Neumann Boundary Conditions
We now turn to the Hölder continuity of weak solutions to equation ($\mathcal{P}$). Since the boundedness condition ($\mathcal{M}$) required in Theorem 1.11 is ensured by Theorems 1.9 and 1.10, we are in a position to establish the Hölder regularity of weak solutions to Problem ($\mathcal{P}$) under both Dirichlet ($\mathcal{PD}$) and Neumann ($\mathcal{PN}$) boundary conditions. This result is summarized in Theorems 1.111.14.
Proof of Theorem 1.11.
Let $\bar{B}_{t}\subset B_{s}\subset\Omega$ and select a $C^{\infty}$-function $\xi$ satisfying
$0\leq\xi\leq 1,\quad\text{supp }\xi\subset B_{s},\quad\xi\equiv 1\text{ on }B_{t},\quad|D\xi|\leq\frac{2}{s-t}.$ (4.29)
Define
$\delta=\min\left\{\frac{a_{4}}{4b_{3}M},2\right\}$ (4.30)
and set $v:=\xi^{g^{+}}(u-k)^{+}=\xi^{g^{+}}\max\{u-k,0\}$, where
$k\geq\max_{B_{s}}u(x)-\delta M.$ (4.31)
Clearly, $v\in W_{0}^{1,G(x,t)}(\Omega)$. Substituting into equation ($\mathcal{FDV}$) yields
$\int_{A_{k,s}}\xi^{g^{+}}A(x,u,Du)\cdot Du\mathrm{d}x$ $+\int_{A_{k,s}}g^{+}\xi^{g^{+}-1}(u-k)A(x,u,Du)\cdot D\xi\mathrm{d}x$ (4.32)
$=\int_{A_{k,s}}B(x,u,Du)\xi^{g^{+}}(u-k)\mathrm{d}x.$ (4.33)
In light of conditions ($\mathcal{A}_{1}$)–($\mathcal{A}_{2}$) and ($\mathcal{B}$), we derive the inequality
$\begin{split}a_{4}\int_{A_{k,s}}G(x,|Du|)\xi^{g^{+}}\mathrm{d}x&\leq a_{5}\int_{A_{k,s}}H_{1}(x,|u|)\xi^{g^{+}}\mathrm{d}x+a_{6}\int_{A_{k,s}}\xi^{g^{+}}\mathrm{d}x\\ &\quad+a_{1}g^{+}\int_{A_{k,s}}g(x,|Du|)\xi^{g^{+}-1}|D\xi|(u-k)\mathrm{d}x\\ &\quad+a_{2}g^{+}\int_{A_{k,s}}P(x,|u|)\xi^{g^{+}-1}|D\xi|(u-k)\mathrm{d}x\\ &\quad+a_{3}g^{+}\int_{A_{k,s}}|D\xi|\xi^{g^{+}-1}(u-k)\mathrm{d}x\\ &\quad+b_{1}\int_{A_{k,s}}M(x,|Du|)\xi^{g^{+}}(u-k)\mathrm{d}x\\ &\quad+b_{2}\int_{A_{k,s}}h_{1}(x,|u|)\xi^{g^{+}}(u-k)\mathrm{d}x\\ &\quad+b_{3}\int_{A_{k,s}}\xi^{g^{+}}(u-k)\mathrm{d}x.\end{split}$ (4.34)
We proceed to estimate each term on the right-hand side of (4.34).
Employing (2.4) and Young’s inequality (2.3), select $\varepsilon_{1}\in(0,1)$ such that
$a_{1}(g^{+})^{2}\varepsilon_{1}^{\frac{g^{+}}{g^{+}-1}}\leq\frac{a_{4}}{4},$
gives that
$\begin{split}&a_{1}g^{+}\int_{A_{k,s}}g(x,|Du|)\xi^{g^{+}-1}|D\xi|(u-k)\mathrm{d}x\\ &\quad\leq a_{1}g^{+}\int_{A_{k,s}}\widetilde{G}\left(x,g(x,|Du|)\xi^{g^{+}-1}\varepsilon_{1}\right)\mathrm{d}x+a_{1}g^{+}\int_{A_{k,s}}G\left(x,\frac{|D\xi|(u-k)}{\varepsilon_{1}}\right)\mathrm{d}x\\ &\quad\leq a_{1}g^{+}\varepsilon_{1}^{\frac{g^{+}}{g^{+}-1}}\int_{A_{k,s}}\widetilde{G}\left(x,g(x,|Du|)\right)\xi^{g^{+}}\mathrm{d}x+a_{1}g^{+}\varepsilon_{1}^{-g^{+}}\int_{A_{k,s}}G\left(x,\frac{2(u-k)}{s-t}\right)\mathrm{d}x\\ &\quad\leq a_{1}(g^{+})^{2}\varepsilon_{1}^{\frac{g^{+}}{g^{+}-1}}\int_{A_{k,s}}G(x,|Du|)\xi^{g^{+}}\mathrm{d}x+a_{1}g^{+}\varepsilon_{1}^{-g^{+}}2^{g^{+}}\int_{A_{k,s}}G\left(x,\frac{u-k}{s-t}\right)\mathrm{d}x\\ &\quad\leq\frac{a_{4}}{4}\int_{A_{k,s}}G(x,|Du|)\xi^{g^{+}}\mathrm{d}x+d_{1}\int_{A_{k,s}}G\left(x,\frac{u-k}{s-t}\right)\mathrm{d}x,\end{split}$ (4.35)
with $d_{1}=d_{1}(a_{1},a_{2},a_{3},g^{-},g^{+})>0$.
An application of Young’s inequality (2.3) and Proposition 2.11, with $\varepsilon_{2}\in(0,1)$ chosen so that
$a_{3}g^{+}2^{g^{+}}\varepsilon_{2}^{g^{-}}\leq 1,$
yields that
$\begin{split}&a_{3}g^{+}\int_{A_{k,s}}|D\xi|(u-k)\xi^{g^{+}-1}\mathrm{d}x\leq a_{3}g^{+}\int_{A_{k,s}}|D\xi|(u-k)\mathrm{d}x\\ &\quad\leq a_{3}g^{+}\int_{A_{k,s}}G\left(x,\varepsilon_{2}|D\xi|(u-k)\right)\mathrm{d}x+a_{3}g^{+}\int_{A_{k,s}}\widetilde{G}\left(x,\frac{1}{\varepsilon_{2}}\right)\mathrm{d}x\\ &\quad\leq a_{3}g^{+}\varepsilon_{2}^{g^{-}}2^{g^{+}}\int_{A_{k,s}}G\left(x,\frac{u-k}{s-t}\right)\mathrm{d}x+a_{3}g^{+}\int_{A_{k,s}}\widetilde{G}\left(x,\frac{1}{\varepsilon_{2}}\right)\mathrm{d}x\\ &\quad\leq\int_{A_{k,s}}G\left(x,\frac{u-k}{s-t}\right)\mathrm{d}x+c_{1}|A_{k,s}|,\end{split}$ (4.36)
where $c_{1}=c_{1}(a_{3},g^{+})>0$.
From ($\mathcal{H}_{1}$), ($\mathcal{M}$), and Proposition 2.11, we find
$\begin{split}a_{5}\int_{A_{k,s}}H_{1}(x,|u|)\xi^{g^{+}}\mathrm{d}x&\leq a_{5}\int_{A_{k,s}}H_{1}(x,M)\xi^{g^{+}}\mathrm{d}x\\ &\leq a_{5}\int_{A_{k,s}}F_{1}M^{h_{1}^{+}}\xi^{g^{+}}\mathrm{d}x\leq c_{2}|A_{k,s}|,\end{split}$ (4.37)
with $c_{2}=c_{2}(a_{5},F_{1},h_{1}^{+},M)>0$.
Invoking Young’s inequality (2.3), Proposition 2.11, inequalities (1.5), (4.37), and selecting $\varepsilon_{3}\in(0,1)$ such that
$a_{2}g^{+}2^{g^{+}}\varepsilon_{3}^{-g^{+}}\leq 1,$
we obtain
$\begin{split}&a_{2}g^{+}\int_{A_{k,s}}P(x,|u|)\xi^{g^{+}-1}|D\xi|(u-k)\mathrm{d}x\\ &\quad\leq a_{2}g^{+}\int_{A_{k,s}}\widetilde{G}\left(x,P(x,|u|)\xi^{g^{+}-1}\varepsilon_{3}\right)\mathrm{d}x+a_{2}g^{+}\int_{A_{k,s}}G\left(x,\frac{|D\xi|(u-k)}{\varepsilon_{3}}\right)\mathrm{d}x\\ &\quad\leq a_{2}g^{+}\varepsilon_{3}^{\frac{g^{+}}{g^{+}-1}}\int_{A_{k,s}}\widetilde{G}\left(x,P(x,|u|)\right)\xi^{g^{+}}\mathrm{d}x+a_{2}g^{+}\varepsilon_{3}^{-g^{+}}\int_{A_{k,s}}G\left(x,\frac{2(u-k)}{s-t}\right)\mathrm{d}x\\ &\quad\leq a_{2}g^{+}\varepsilon_{3}^{\frac{g^{+}}{g^{+}-1}}\int_{A_{k,s}}H_{1}(x,|u|)\xi^{g^{+}}\mathrm{d}x+a_{2}g^{+}\varepsilon_{3}^{-g^{+}}2^{g^{+}}\int_{A_{k,s}}G\left(x,\frac{u-k}{s-t}\right)\mathrm{d}x\\ &\quad\leq c_{3}|A_{k,s}|+d_{2}\int_{A_{k,s}}G\left(x,\frac{u-k}{s-t}\right)\mathrm{d}x,\end{split}$ (4.38)
where $c_{3}=c_{3}(a_{2},F_{1},h_{1}^{+},M,g^{+})>0$ and $d_{2}=d_{2}(a_{2},g^{+})>0$.
In light of (1.4), ($\mathcal{H}_{1}$), ($\mathcal{M}$), Proposition 2.11, and Young’s inequality (2.3), choose $\varepsilon_{4}\in(0,1)$ such that
$b_{1}\varepsilon_{4}^{\frac{h^{+}_{1}}{h^{+}_{1}-1}}\leq\frac{a_{4}}{4},$
we see that
$\begin{split}&b_{1}\int_{A_{k,s}}M(x,|Du|)\xi^{g^{+}}(u-k)\mathrm{d}x\\ &\quad\leq b_{1}\int_{A_{k,s}}\widetilde{H_{1}}\left(x,\varepsilon_{4}M(x,|Du|)\xi^{g^{+}}\right)\mathrm{d}x+b_{1}\int_{A_{k,s}}H_{1}\left(x,\frac{u-k}{\varepsilon_{4}}\right)\mathrm{d}x\\ &\quad\leq b_{1}\varepsilon_{4}^{\frac{h^{+}_{1}}{h^{+}_{1}-1}}\int_{A_{k,s}}\widetilde{H_{1}}\left(x,M(x,|Du|)\right)\xi^{g^{+}\frac{h^{+}_{1}}{h^{+}_{1}-1}}\mathrm{d}x+b_{1}\int_{A_{k,s}}H_{1}\left(x,\frac{\delta M}{\varepsilon_{4}}\right)\mathrm{d}x\\ &\quad\leq b_{1}\varepsilon_{4}^{\frac{h^{+}_{1}}{h^{+}_{1}-1}}\int_{A_{k,s}}G(x,|Du|)\xi^{g^{+}}\mathrm{d}x+b_{1}(\delta M)^{h_{1}^{+}}\varepsilon_{4}^{-h^{+}_{1}}F_{1}|A_{k,s}|\\ &\quad\leq\frac{a_{4}}{4}\int_{A_{k,s}}G(x,|Du|)\xi^{g^{+}}\mathrm{d}x+c_{4}|A_{k,s}|,\end{split}$ (4.39)
with $c_{4}=c_{4}(b_{1},M,\delta,F_{1},h_{1}^{+})>0$.
From (2.4), ($\mathcal{H}_{1}$), ($\mathcal{M}$), Proposition 2.11, and Young’s inequality (2.3), we get
$\begin{split}&b_{2}\int_{A_{k,s}}h_{1}(x,|u|)\xi^{g^{+}}(u-k)\mathrm{d}x\\ &\quad\leq b_{2}\int_{A_{k,s}}\widetilde{H_{1}}\left(x,h_{1}(x,|u|)\right)\mathrm{d}x+b_{2}\int_{A_{k,s}}H_{1}(x,(u-k))\mathrm{d}x\\ &\quad\leq b_{2}h^{+}_{1}\int_{A_{k,s}}H_{1}\left(x,|u|\right)\mathrm{d}x+b_{2}\int_{A_{k,s}}H_{1}(x,\delta M)\mathrm{d}x\\ &\quad\leq b_{2}h^{+}_{1}F_{1}M^{h_{1}^{+}}|A_{k,s}|+b_{2}F_{1}(\delta M)^{h^{+}_{1}}|A_{k,s}|\\ &\quad\leq\left(b_{2}F_{1}M^{h_{1}^{+}}\right)\left(\delta^{h^{+}_{1}}+h_{1}^{+}\right)|A_{k,s}|\leq c_{5}|A_{k,s}|,\end{split}$ (4.40)
where $c_{5}=c_{5}(b_{2},M,\delta,F_{1},h_{1}^{+})>0$.
The remaining terms are bounded directly. From (4.30) and (4.31), we have
$a_{6}\int_{A_{k,s}}\xi^{g^{+}}\mathrm{d}x$ $\leq a_{6}|A_{k,s}|,$ (4.41)
$\int_{A_{k,s}}\xi^{g^{+}}(u-k)\mathrm{d}x$ $\leq\int_{A_{k,s}}(u-k)\mathrm{d}x\leq(\max_{B_{s}}u(x)-k)|A_{k,s}|\leq\delta M|A_{k,s}|\leq\frac{a_{4}}{4b_{3}}|A_{k,s}|.$ (4.42)
Combining estimates (4.34)–(4.42), we conclude that
$\int_{A_{k,s}}G(x,|Du|)\xi^{g^{+}}\mathrm{d}x\leq\gamma\int_{A_{k,s}}G\left(x,\frac{u-k}{s-t}\right)\mathrm{d}x+\gamma_{1}|A_{k,s}|,$
and consequently,
$\int_{A_{k,t}}G(x,|Du|)\mathrm{d}x\leq\gamma\int_{A_{k,s}}G\left(x,\frac{u-k}{s-t}\right)\mathrm{d}x+\gamma_{1}|A_{k,s}|,$ (4.43)
for $\bar{B}_{t}\subset B_{s}\subset\Omega$ and $k$ satisfying (4.31), where
$\gamma:=\gamma(a_{2},a_{3},a_{4},g^{-},g^{+})\quad\text{and}\quad\gamma_{1}:=\gamma_{1}(a_{2},a_{3},a_{5},a_{6},b_{1},b_{2},b_{3},g^{+},h^{+}_{1},F_{1},M)$
are positive constants. Inequality (4.43) implies $u\in\mathcal{B}_{G(x,t)}(\Omega,M,\gamma,\gamma_{1},\delta)$. A similar argument shows
$u\in\mathcal{B}_{G(x,t)}(\overline{\Omega},M,\gamma,\gamma_{1},\delta),$
which completes the proof of Theorem 1.11. ∎
Proof of Theorem 1.12.
The proof proceeds along the same lines as that of Theorem 1.11, and we therefore omit the details. The only modification arises in the treatment of inequality (4.34). Specifically, by using equation ($\mathcal{FNV}$) together with the structural conditions ($\mathcal{A}_{1}$)–($\mathcal{A}_{2}$), ($\mathcal{B}$) and ($\mathcal{C}$), we obtain the following estimate
$\begin{split}a_{4}\int_{A_{k,s}}G(x,|Du|)\xi^{g^{+}}{\rm d}x&\leq a_{5}\int_{A_{k,s}}H_{1}(x,|u|)\xi^{g^{+}}{\rm d}x+a_{6}\int_{A_{k,s}}\xi^{g^{+}}{\rm d}x+a_{1}g^{+}\int_{A_{k,s}}g(x,|Du|)\xi^{g^{+}-1}|D\xi|(u-k){\rm d}x\\ &+a_{2}g^{+}\int_{A_{k,s}}P(x,|u|)\xi^{g^{+}-1}|D\xi|(u-k){\rm d}x+a_{3}g^{+}\int_{A_{k,s}}|D\xi|\xi^{g^{+}-1}(u-k){\rm d}x\\ &+b_{1}\int_{A_{k,s}}M(x,|Du|)\xi^{g^{+}}(u-k){\rm d}x+b_{2}\int_{A_{k,s}}h_{1}(x,|u|)\xi^{g^{+}}(u-k){\rm d}x\\ &+b_{3}\int_{A_{k,s}}\xi^{g^{+}}(u-k){\rm d}x+c_{1}\int_{\Gamma_{k,s}}h_{2}(x,|u|)\xi^{g^{+}}(u-k){\rm d}x\\ &+c_{2}\int_{\Gamma_{k,s}}\xi^{g^{+}}(u-k){\rm d}x\\ &\leq a_{5}\int_{A_{k,s}}H_{1}(x,|u|)\xi^{g^{+}}{\rm d}x+a_{6}\int_{A_{k,s}}\xi^{g^{+}}{\rm d}x+a_{1}g^{+}\int_{A_{k,s}}g(x,|Du|)\xi^{g^{+}-1}|D\xi|(u-k){\rm d}x\\ &+a_{2}g^{+}\int_{A_{k,s}}P(x,|u|)\xi^{g^{+}-1}|D\xi|(u-k){\rm d}x+a_{3}g^{+}\int_{A_{k,s}}|D\xi|\xi^{g^{+}-1}(u-k){\rm d}x\\ &+b_{1}\int_{A_{k,s}}M(x,|Du|)\xi^{g^{+}}(u-k){\rm d}x+b_{2}\int_{A_{k,s}}h_{1}(x,|u|)\xi^{g^{+}}(u-k){\rm d}x\\ &+(b_{3}+c_{2})\int_{A_{k,s}}\xi^{g^{+}}(u-k){\rm d}x+c_{1}\int_{A_{k,s}}h_{2}(x,|u|)\xi^{g^{+}}(u-k){\rm d}x,\end{split}$ (4.44)
where, $\Gamma_{k,s}:=\left\{x\in B_{s}\cap\partial\Omega:u(x)>k\right\}.$
Proof of Theorems 1.13 and 1.14.
Theorems 1.13 and 1.14 follow as direct consequences of Theorems 1.21.12. ∎
Author Contributions: The authors contributed equally to this work.
Data Availability: Data sharing is not applicable to this article as no new data were created or analyzed in this study.
Declarations
Ethical Approval: Not applicable.
Conflict of interest: The authors declare that they have no Conflict of interest.